### v2

# Byzantine Failures Harm the Generalization of Robust Distributed Learning Algorithms More Than Data Poisoning

Abstract
Robust distributed learning algorithms aim to maintain reliable performance despite the presence of misbehaving workers. Such misbehaviors are commonly modeled as Byzantine failures, allowing arbitrarily corrupted communication, or as data poisoning, a weaker form of corruption restricted to local training data. While prior work shows similar optimization guarantees for both models, an important question remains: How do these threat models impact generalization? Empirical evidence suggests a gap, yet it remains unclear whether it is unavoidable or merely an artifact of suboptimal attacks. We show, for the first time, a fundamental gap in generalization guarantees between the two threat models: Byzantine failures yield strictly worse rates than those achievable under data poisoning. Our findings leverage a tight algorithmic stability analysis of robust distributed learning. Specifically, we prove that: (i) under data poisoning, the uniform algorithmic stability of an algorithm with optimal optimization guarantees degrades by an additive factor of $\varTheta(\frac{f}{n-f})$, with $f$ out of $n$ workers misbehaving; whereas (ii) under Byzantine failures, the degradation is in $\Omega\big(\sqrt{\frac{f}{n-2f}}\big)$.
1 INTRODUCTION
With the proliferation of large-scale data and pervasive connectivity, distributed learning has evolved into a fundamental paradigm in modern AI (Kairouz et al., 2021). From federated smartphones to geo-distributed datacenters, algorithms now orchestrate millions of workers toward shared objectives. Yet, unlike the idealized setting of perfectly cooperative workers, real-world systems face a range of adversarial misbehaviors, broadly categorized under the umbrella of Byzantine failures (Guerraoui et al., 2024). Encompassing everything from intermittent data corruption to sophisticated, coordinated attacks, these failures threaten the reliability of distributed learning algorithms.
Robust distributed learning aims to maintain strong learning guarantees despite the presence of such misbehaviors. The prevailing approach is to model a fraction of workers as corrupted, either Byzantine—able to send arbitrarily corrupted updates—or poisoned, where corruption is limited to local training data. While Farhadkhani et al. (2024b) showed that both threat models yield similar empirical risk guarantees, experiments consistently observe significantly worse generalization—i.e., higher risk on unseen data—under Byzantine failures (Allouah et al., 2023a; Karimireddy et al., 2022). Whether this gap reflects an inherent difference in adversarial power or only the suboptimality of practical attacks remains unknown. We address this open problem with the first theoretical analysis of the generalization gap between Byzantine failures and data poisoning.
To this end, we use algorithmic stability (Bousquet and Elisseeff, 2002) to quantify generalization resilience against both forms of attacks. In particular, uniform stability—which measures an algorithm’s sensitivity to the replacement of a single training example—offers a worst-case, distribution-free, and algorithm-centric perspective on generalization. This stands in contrast to approaches based on uniform convergence (Vapnik, 1998), PAC-Bayes (McAllester, 1999), or mutual information (Russo and Zou, 2019), which provide less direct insight into algorithm-specific behavior. This makes stability well-suited for addressing the open question under consideration, as it enables a focused analysis of the potential harm posed by different threat models on the generalization performance of robust learning algorithms.
Our contributions.
[…]
2 PRELIMINARIES
Problem setting.
We consider a distributed setup with a central server and $n$ workers, where up to $f$ (with $f<\frac{n}{2}$) may be subject to either Byzantine failures or data poisoning, as defined below. We refer to such workers as misbehaving, unless we need to distinguish between the two attack types. Their identities are unknown to the server. While the actual number of misbehaving workers may be less than $f$, we assume the worst-case scenario where exactly $f$ workers are misbehaving. The remaining honest workers form the set $\mathcal{H}\subseteq[n]\vcentcolon=\{1,\ldots,n\}$, with $|\mathcal{H}|=n-f$. Each honest worker $i\in\mathcal{H}$ holds a local dataset $\mathcal{D}_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}$ composed of $m$ i.i.d. data points (or samples) from an input space $\mathcal{Z}$ drawn from a distribution $p_{i}$. We denote $\mathcal{S}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Given a parameter vector $\theta\in\Theta\subset\mathbb{R}^{d}$ representing the model, a data point $z\in\mathcal{Z}$ incurs a differentiable loss defined by a real-valued function $\ell(\theta;z)$. The goal is to minimize the population risk over the honest workers:
$R_{\mathcal{H}}(\theta)=\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\mathbb{E}_{z\sim p_{i}}[\ell(\theta,z)].$
As we only have access to a finite number of samples from each distribution, the above objective can only be solved approximately by minimizing the empirical risk over the honest workers, with the added challenge that the honest subgroup is unknown:
$\widehat{R}_{\mathcal{H}}(\theta)=\frac{1}{|\mathcal{H}|}\!\sum_{i\in\mathcal{H}}\widehat{R}_{i}(\theta)=\frac{1}{|\mathcal{H}|}\!\sum_{i\in\mathcal{H}}\frac{1}{m}\!\sum_{z\in\mathcal{D}_{i}}\ell(\theta,z).$
The expected excess population risk of a distributed learning algorithm’s output $\mathcal{A(\mathcal{S})}$ can be decomposed into the generalization and optimization errors (see, e.g. Hardt et al. (2016)) as follows:
$\mathbb{E}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\inf\limits_{\theta\in\Theta}R_{\mathcal{H}}(\theta)\big]\leq\mathbb{E}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\\ -\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]+\mathbb{E}\big[\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\inf\limits_{\theta\in\Theta}\widehat{R}_{\mathcal{H}}(\theta)\big]$ (1)
Prior work on robust distributed learning has focused primarily on the optimization error, i.e., the second term on the right-hand side. In contrast, our goal is to study the generalization error, corresponding to the first term. Formally, we define the notion of robustness as follows, with $\|\cdot\|_{2}$ the Euclidean norm.
Definition 2.1.
An algorithm is said to be $(f,\rho,$ statistical)-resilient if it outputs a parameter $\hat{\theta}$ s. t.
$\mathbb{E}_{\mathcal{A}}\big[R_{\mathcal{H}}(\hat{\theta})\!-\!\inf_{\theta\in\Theta}R_{\mathcal{H}}(\theta)\big]\!\leq\!\rho.$
We say $(f,\rho,\text{empirical})$-resilient when we only consider empirical risk.
Note that the (standard) notion of $(f,\rho,\text{empirical})$-resilience—see, e.g., Liu et al. (2021)—is impossible in general for any $\rho$ when $f\geq n/2$, justifying our assumption $f<n/2$.
Threat models. We aim to analyze this resilience property under two standard threat models widely studied in the literature (Farhadkhani et al., 2024a; Shejwalkar and Houmansadr, 2021).
Byzantine failures. In this threat model, misbehaving workers can act arbitrarily: they may deviate from the prescribed algorithm, collude, and send adversarial updates to the server while having access to all information exchanged between honest workers and the server. This model follows the classical notion of Byzantine failures in distributed systems introduced in the seminal work of Lamport et al. (1982).
Data poisoning. In this restricted threat model, misbehaving workers follow the prescribed algorithm but may corrupt their local training data before execution. Specifically, for each misbehaving worker $i\notin\mathcal{H}$, a local dataset $\mathcal{D}_{i}\in\mathcal{Z}^{m}$ is adversarially constructed with full knowledge of the honest workers’ data distributions $\{p_{i},~i\in\mathcal{H}\}$.
Background on robust distributed optimization.
Robust distributed optimization algorithms aim to achieve $(f,\rho,\text{empirical})$-resilience. They are typically adaptations of standard first-order iterative optimization algorithms like gradient descent ($\mathrm{GD}$) or stochastic gradient descent ($\mathrm{SGD}$). These algorithms are made resilient by replacing the server-side averaging operator by a robust aggregation rule $F:(\mathbb{R}^{d})^{n}\to\mathbb{R}^{d}$ (Guerraoui et al., 2024). Formally, given a learning rate $\gamma$, the parameter $\theta_{t}$ at iteration $t\in\{0,\ldots,T-1\}$ is updated as follows:
$\theta_{t+1}=G^{F}_{\gamma}(\theta_{t}):=\theta_{t}-\gamma F(g^{(1)}_{t},\ldots,g^{(n)}_{t})$,
where $g^{(i)}_{t}$ denotes the update sent by worker $i$ at iteration $t$. For an honest worker $i\in\mathcal{H}$, we have $g^{(i)}_{t}=\nabla\widehat{R}_{i}(\theta_{t})$ under $\mathrm{GD}$, or $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i)}_{t})$ under $\mathrm{SGD}$, with $z^{(i)}_{t}$ uniformly sampled from the local dataset $\mathcal{D}_{i}$. For a misbehaving worker $i\notin\mathcal{H}$, $g^{(i)}_{t}$ is either an arbitrary vector in $\mathbb{R}^{d}$ (Byzantine failures) or a gradient computed on corrupted data (data poisoning).
Recent advances in robust distributed optimization have identified key properties sufficient for a robust aggregation rule to ensure $(f,\rho,\text{empirical})$-resilience (Allouah et al., 2023a, b). The following definition encompasses a wide range of aggregation rules and has been shown to yield tight resilience guarantees across a variety of practical distributed learning settings.
Definition 2.2.
Let $n\!\geq\!1$, $0\!\leq f\!<\!n/2$ and $\kappa\!\geq\!0$. An aggregation rule $F$ is $(f,\kappa)$-robust if for any $g_{1},\ldots,g_{n}\in\mathbb{R}^{d}$, any set $S\!\subset\![n]$ of size nfn-f, with $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g_{i}$ and $\Sigma_{S}=\frac{1}{|S|}\sum_{i\in S}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal}$,
$\|F(g_{1},\ldots,g_{n})-\overline{g}_{S}\|_{2}^{2}\leq\kappa\|\Sigma_{S}\|_{sp}.$
Here, $\|\cdot\|_{sp}$ denotes the spectral norm (i.e., the largest eigenvalue), $f$ and $\kappa$ are referred to as the robustness parameter and robustness coefficient of $F$, respectively.
Any aggregation rule $F$ that is $(f,\mathcal{O}(f/n))$-robust achieves optimal $(f,\rho,\text{empirical})$-resilience. In particular, iterative methods employing such a rule attain an optimization error that matches the information-theoretic lower bound for first-order optimization in the presence of $f$ Byzantine workers. One such optimally robust rule is SMEA (Allouah et al., 2023b).
Definition 2.3.
Given $f,n\in\mathbb{N}$, $f<n/2$ and vectors $g_{1},\ldots,g_{n}\in\mathbb{R}^{d}$, the smallest maximum eigenvalue averaging aggregation rule (SMEA) outputs the average of the nfn-f gradients that are most directionally consistent, discarding outliers based on covariance:
$\overline{g}_{S^{*}}=\frac{1}{|S^{*}|}\sum_{i\in S^{*}_{t}}g_{i},\quad\text{with}\quad S^{*}\in\underset{\begin{subarray}{c}{S\subseteq\{1,\ldots,n\},}\\ {|S|=n-f}\end{subarray}}{\arg\!\min}\|\Sigma_{S}\|_{sp}$
However, the above results focus on the optimization error, overlooking the impact of robust aggregation rules on the generalization error—the first term in the right-hand side of eq. 1. To address this gap, we provide a rigorous analysis of the generalization guarantees of robust distributed learning, focusing on the $\mathrm{SMEA}$ aggregation rule as it yields optimal optimization guarantees.
Stability under misbehaving workers.
Our study leverages the framework of algorithmic stability to understand the generalization error of robust distributed learning algorithms. We do so via uniform algorithmic stability, which measures how sensitive a learning algorithm is to changes in its training data. This approach was introduced by Vapnik and Chervonenkis (1974); Devroye and Wagner (1979); Rogers and Wagner (1978) and later popularized by Bousquet and Elisseeff (2002); Shalev-Shwartz et al. (2010); Hardt et al. (2016). We revisit this framework to study generalization in robust distributed learning, where the change is only in the honest workers’ data.
Definition 2.4.
Consider $n$ workers, with $f$ misbehaving and nfn-f honest, each holding $m$ local samples. A distributed algorithm $\mathcal{A}$ is $\varepsilon$-uniformly stable if, for all honest datasets $\mathcal{S},\mathcal{S}^{\prime}\in\mathcal{Z}^{(n-f)m}$ that are neighboring—i.e., differing in at most one sample—we have
$\sup_{z\in\mathcal{Z}}\mathbb{E}[\ell(\mathcal{A}(\mathcal{S});z)-\ell(\mathcal{A}(\mathcal{S}^{\prime});z)]\leq\varepsilon,$
where the expectation is over the randomness of $\mathcal{A}$.
This property leads to the following bound on the generalization error, proved in Appendix A.
Proposition 2.1.
If $\mathcal{A}$ is $\varepsilon$-uniformly stable, then
$|\mathbb{E}_{\mathcal{S},\mathcal{A}}[R_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))-\widehat{R}_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))]|\leq\varepsilon.$
The link between uniform stability and generalization can be further strengthened to yield high-probability generalization bounds, as discussed in Appendix G. We can now formalize a relationship between empirical resilience and statistical resilience through stability: any algorithm that is $(f,\rho,\text{empirical})$-resilient and $\varepsilon$-uniformly stable is $(f,\rho+\varepsilon,\text{statistical})$-resilient. Here, $\varepsilon$ quantifies the effect of misbehaving workers on generalization, and is the central focus of this paper.
3 (IN)STABILITY OF ROBUST DISTRIBUTED LEARNING
In this section, we first derive upper bounds on the stability of robust distributed $(\mathrm{S})\mathrm{GD}$ under general $(f,\kappa)$-robust aggregation rules. We next analyze $\mathrm{SMEA}$ to derive lower bounds under Byzantine failures, as well as tight upper and lower bounds for the special case of data poisoning. Our results show a fundamental gap in stability between the two threat models. Definitions of function regularity are given in Appendix B.
3.1 The Case of Byzantine Failures
Upper bounds. We start by deriving uniform stability upper bounds for convex optimization. Proofs are deferred to Section C.2.
Theorem 3.1.
Consider the setting described in Section 2 under Byzantine failures. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ with a $(f,\kappa)$-robust aggregation rule $F$. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth, and $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$.
  1. (i)
    If $\ell(\cdot;z)$ is convex $\forall z\in\mathcal{Z}$, then the uniform stability of $\mathcal{A}$ is upper bounded by
    $2\gamma C^{2}T\left(\sqrt{\kappa}+\frac{1}{(n-f)m}\right).$ (2)
  2. (ii)
    If $\ell(\cdot;z)$ is $\mu$-strongly convex $\forall z\in\mathcal{Z}$, then the uniform stability of $\mathcal{A}$ is upper bounded by
    $\frac{2C^{2}}{\mu}\left(\sqrt{\kappa}+\frac{1}{(n-f)m}\right).$ (3)
When there are no Byzantine workers ($f=0$, hence $\kappa=0$), both eqs. 2 and 3 reduce to the classical stability bounds established in Hardt et al. (2016). However, when $f>0$, the analysis reveals a degradation by an additive term of order $\mathcal{O}\left(\sqrt{\kappa}\right)$. In the convex setting, this degradation accumulates over the $T$ iterations, whereas in the strongly convex case, it appears as a one-time term independent of $T$.
Lower bounds. Note that $\kappa\geq\frac{f}{n-2f}$ (Allouah et al., 2023a, Proposition 6). Therefore, the additive term in Theorem 3.1 is lower bounded by $\sqrt{\frac{f}{n-2f}}$. We show that this bound is tight by deriving the following lower bound, proved in Section C.3.
Theorem 3.2.
Consider the setting in Section 2 under Byzantine failures and assume $\frac{n}{3}\!\leq\!f\!<\!\frac{n}{2}$. Let $\mathcal{A}=\mathrm{GD}$, with $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\!\in\!\mathbb{N}^{*}$ iterations with learning rate $\gamma$. Then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and neighboring datasets such that the uniform stability of $\mathcal{A}$ is lower bounded by
$\Omega\left(\gamma C^{2}T\left(\sqrt{\frac{f}{n-2f}}+1+\frac{1}{(n-f)m}\right)\right).$ (4)
Proof sketch..
Let $C,L\in\mathbb{R}_{+}^{*}$ and let $\ell(\theta;z)=z\theta$, $\theta\in\mathbb{R}$, $z\in[-C,C]$, which is $C$-Lipschitz, $L$-smooth and convex. In this particular case, the $\mathrm{SMEA}$ rule amounts to picking the subset of size nfn-f with smallest variance. Our proof relies on three key elements. (i) We maximize the variance among honest gradients by dividing them into two equal-sized subgroups positioned at the boundary of the Lipschitz constraint. More precisely, $\frac{n-f}{2}$ honest workers have datasets with $m$ identical samples equal to $C$, the remaining $\frac{n-f}{2}$ have samples equal to C-C. Hence, this set of nfn-f values has, at each iteration, a variance equal to $C^{2}$. (ii) When $f\geq\frac{n}{3}$, Byzantine workers can craft updates that cause $\mathrm{SMEA}$ to entirely discard one honest subgroup while outweighing the influence of the other. Specifically, if Byzantine workers send $\beta=C\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}\right)$ at each iteration, then the subgroup consisting of $f\geq\frac{n-f}{2}$ Byzantine workers together with $n-2f\leq\frac{n-f}{2}$ honest workers sending $C$ forms the subset with the lowest variance, equal to $C^{2}/4<C^{2}$. Consequently, under this configuration, the parameter at step $T$ becomes $\theta_{T}=-\gamma CT\left(1+\frac{1}{2}\sqrt{\frac{f}{n-2f}}\right)$. (iii) Finally, we exploit the fact that Byzantine workers can observe all communications and adapt their behavior accordingly. Specifically, we can define an event that triggers them to switch their communicated value $\beta$ to an arbitrarily large number. For instance, if a single sample from a honest worker is changed from $C$ to 0—defining a neighboring dataset—the server receives a different average $\frac{m-1}{m}C<C$ that triggers the Byzantine workers. The choice of an arbitrary large value is a matter of convenience, ensuring that the perturbed run diverges in the positive direction at a rate of $\theta_{T}^{\prime}=\frac{\gamma CT}{(n-f)m}$. As, in this setting, stability is proportional to $\left|\theta_{T}-\theta_{T}^{\prime}\right|$, this concludes the proof sketch.∎
On the tightness of our bounds.
Note that, for $\mathrm{SMEA}$, $\sqrt{\kappa}\in\mathcal{O}\Big(\sqrt{\frac{f}{n-f}}\big(1+\frac{f}{n-2f}\big)\Big)$ (Allouah et al., 2023b, Proposition 5.1). Plugging this in our upper-bound and considering the regime $n/3\!\leq f\!\leq\!n/(2+\nu)$ for some constant $\nu>0$, our upper and lower bounds become tight at a rate $\varTheta\Big(\sqrt{\frac{f}{n-f}}\Big)$, providing a precise characterization of stability under Byzantine failures. For $f<n/3$, we lack a strong lower bound; nonetheless, we investigate this regime numerically in Section 4, yielding valuable insights into the transition from few to many misbehaving workers.
Crucially, the above lower bound cannot be reproduced under data poisoning, due to points (ii) and (iii) in the proof sketch. Unlike Byzantine workers, who can send arbitrary updates adaptively, poisoned workers are constrained by the loss function’s regularity and by being committed to a fixed dataset before the algorithm begins. The next section formally establishes a stability gap between the two threat models.
3.2 The Case of Data Poisoning
Unlike our previous bounds, which account for the arbitrariness of Byzantine vectors, we now leverage the regularity of poisoned gradients in data poisoning to derive tighter stability bounds for robust distributed learning with $\mathrm{SMEA}$. Moreover, as outlined in the proof sketch below, the core reasoning underlying our analysis extends more broadly: it provides a general approach for lower bounding uniform stability under $(f,\kappa)$-robust aggregation rules that do not preserve the co-coercivity of the loss function (see eq. 13 in Appendix B). Proofs are in Sections D.1, D.2 and D.3.
Upper bounds. We start by proving a tighter upper bound under data poisoning.
Theorem 3.3.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with $\mathrm{SMEA}$. Suppose $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth $\forall z\!\in\mathcal{Z}$, and $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$.
  1. (i)
    If $\ell(\cdot;z)$ is convex $\forall z\in\mathcal{Z}$, then the uniform stability of $\mathcal{A}$ is upper bounded by
    $2\gamma C^{2}T\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right).$ (5)
  2. (ii)
    If $\ell(\cdot;z)$ is $\mu$-strongly convex $\forall z\in\mathcal{Z}$, then the uniform stability of $\mathcal{A}$ is upper bounded by
    $\frac{2C^{2}}{\mu}\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right).$ (6)
This result generalizes beyond $\mathrm{SMEA}$, yielding the same upper bounds for any rule that averages nfn-f gradients, regardless of the selection criterion.
Lower bounds. Before comparing the above bounds to the Byzantine case, we establish their tightness by deriving matching lower bounds. These lower bounds are established for $\mathrm{projected}\text{--}\mathrm{SGD}$, which applies a Euclidean projection onto the positive half parameter space after each iteration. While this simplifies the lower bounds proofs (see Section D.3), it does not affect the tightness of our result since the upper bounds on uniform stability remain valid for $\mathrm{projected}\text{--}\mathrm{SGD}$ as the projection does not increase the distance between projected points.
Theorem 3.4.
Consider the setting of Section 2 under data poisoning. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{projected}\text{--}\mathrm{SGD}\}$, with $\mathrm{SMEA}$, run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$.
  1. (i)
    There exists $\ell$ s.t. $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and neighboring datasets s.t. the uniform stability of $\mathcal{A}$ is lower bounded by
    $\Omega\left(\gamma C^{2}T\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right)\right).$ (7)
    $\mathcal{A}=\mathrm{projected}\text{--}\mathrm{SGD}$ assumes there exists a constant $\tau\geq c>0$ (arbitrary $c$) such that $T\geq\tau m$.
  2. (ii)
    For $\mathcal{A}\!=\!\mathrm{GD}$, if $T\!\geq\!\frac{\ln(1-c)}{\ln(1-\gamma\mu)}$ with 0<c<10\!<\!c\!<\!1, there exist $\ell$ s.t. $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is additionally $\mu$-strongly convex, and neighboring datasets s.t. the uniform stability of $\mathcal{A}$ is lower bounded by
    $\Omega\left(\frac{C^{2}}{\mu}\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right)\right).$ (8)
Proof sketch..
For formal proofs, see Sections D.2 and D.3. We provide a general rationale for aggregation rules that fail to preserve the co-coercive inequality of the underlying loss. In particular, as detailed below, an additive term necessarily arises in worst-case updates, and this behavior is expected to extend to iterative algorithms. Let $F$ be a robust aggregation rule and $\ell$ a smooth and convex loss function. Suppose there exist $\theta,\omega\!\in\!\Theta$, $\theta\neq\omega$ such that $F$ breaks the co-coercive inequality (eq. 13), that is, denoting $F_{\theta}=F(\nabla\widehat{R}_{1}(\theta),\ldots,\nabla\widehat{R}_{n}(\theta))$ and $F_{\omega}=F(\nabla\widehat{R}_{1}(\omega),\ldots,\nabla\widehat{R}_{n}(\omega))$:
$\langle\theta-\omega,F_{\theta}-F_{\omega}\rangle\leq 0.$ (9)
Equation 9 implies:
$\|G^{F}_{\gamma}(\theta)-G^{F}_{\gamma}(\omega)\|_{2}^{2}=\|\theta-\omega\|_{2}^{2}+\gamma^{2}\|F_{\theta}-F_{\omega}\|_{2}^{2}\\ -2\gamma\langle\theta-\omega,F_{\theta}-F_{\omega}\rangle>\|\theta-\omega\|_{2}^{2},$
Specifically, the quantity ${\sigma^{F}_{\theta,\omega}}=\|G^{F}_{\gamma}(\theta)-G^{F}_{\gamma}(\omega)\|_{2}-\|\theta-\omega\|_{2}>0$ captures the additive growth in distance between parameters due to the update. To illustrate the behavior of $\sigma^{F}_{\theta,\omega}$, we examine representative cases under aggregation rules such as $\mathrm{SMEA}$, which average gradients over an adaptively selected subset. This selection of workers can violate the loss function’s co-coercivity inequality: let $S_{\theta}$ and $S_{\omega}$ denote the subsets selected for parameters $\theta$ and $\omega$ respectively; violation occurs when $S_{\theta}\neq S_{\omega}$. The intersection and difference of the two subsets satisfy: $n-2f\leq|S_{\theta}\cap S_{\omega}|\leq n-f\quad\text{and}\quad|S_{\theta}\setminus S_{\omega}\cup S_{\omega}\setminus S_{\theta}|\leq 2f.$ Hence, through straightforward algebraic manipulation and leveraging gradient boundedness, we obtain in the worst case scenario ${\sigma^{F}_{\theta,\omega}}\propto\frac{f}{n-f}$.
We specifically sketch below our construction for $\mathcal{A}=\mathrm{GD}$, which will be used extensively in Section 4. In dimension one, the $\mathrm{SMEA}$ rule amounts to picking the subset with the smallest gradient variance. For $n=3$ and $f=1$, this reduces to minimizing in $i,j$, $\|g_{i}-g_{j}\|_{2}^{2}/4$, so $\mathrm{SMEA}$ averages the closest pair of worker gradients. Let $\ell(\theta;z)=z\theta$, $\theta\in\mathbb{R}$, $z\in[-C,C]$, and set local datasets to be homogeneous, equal to 0 (worker 1), C-C (worker 2), and $C-\delta$ (worker 3, $\delta>0$). Then $\nabla\widehat{R}_{1}(\theta)=0$, $\nabla\widehat{R}_{2}(\theta)=-C$, $\nabla\widehat{R}_{3}(\theta)=C-\delta$, so $\mathrm{SMEA}$ favors $\{1,3\}$, updating the parameter in the negative direction. Since $C-\delta$ can be made arbitrarily close to $C$, a perturbation of worker 1’s dataset (e.g., changing a sample from 0 to C-C, giving $\nabla\widehat{R}^{\prime}_{1}(\theta)=-\frac{C}{m}$) makes $\mathrm{SMEA}$ favor $\{1,2\}$, updating the parameter in the positive direction. Hence, parameter divergence is proportional to $\gamma CT$. This reasoning generalizes to any $f$ and $n$ (see Section D.2.1), replacing a single sample can cause $\mathrm{SMEA}$ to swap $f$ workers among the nfn-f originally subsampled, yielding stability proportional to $\gamma C^{2}T\Big(\frac{f}{n-f}+\frac{1}{(n-f)m}\Big)$, where the first term reflects worker swapping and the second the perturbed sample. ∎
The failure of robust aggregation rules to preserve the loss function’s co-coercivity inequality underlies the additional instability observed in eq. 7, compared to the simple averaging rule. To our knowledge, no robust aggregation rule preserves this inequality. Identifying such a rule—or proving its impossibility—remains an important open question for future work.
Fundamental gap with Byzantine bounds.
We now compare the bounds obtained for the two threat models. Crucially, the additional instability term in the Byzantine upper bounds (eq. 2 and 3) decreases from $\mathcal{O}(\sqrt{\kappa})$—which for $\mathrm{SMEA}$ evaluates to $\mathcal{O}\big(\sqrt{\frac{f}{n-f}}(1\!+\!\frac{f}{n-2f})\big)$ (Allouah et al., 2023b, Proposition 5.1)—down to $\mathcal{O}(\frac{f}{n-f})$ in the data poisoning case (eq. 5 and 6). Importantly, the upper bound in eq. 5 is thus an order of magnitude smaller than the lower bound $\Omega(\sqrt{\frac{f}{n-2f}})$ eq. 4 obtained for $f\geq\frac{n}{3}$, establishing a fundamental stability gap. As $f$ approaches its maximal value $\tfrac{n}{2}$, the gap widens significantly, highlighting the stability contrast between the two threat models. Table 1 summarizes our results.
Regime Byzantine failures Data poisoning
$f<\frac{n}{3}$ $\mathcal{O}(\sqrt{\frac{f}{n-f}})$ $\varTheta(\frac{f}{n-f})$
$\frac{n}{3}\leq f<\frac{n}{2+\nu}$ $\varTheta(\sqrt{\frac{f}{n-f}})$ $\varTheta(\frac{f}{n-f})$
$f\sim\frac{n}{2}$ $\Omega(\sqrt{\frac{f}{n-2f}})$ $\varTheta(\frac{f}{n-f})$
Table 1: Uniform stability in different threat models.
In the next section, we show that this stability gap translates into a gap in generalization error, revealing a fundamental difference in the ability of the two threat models to degrade performance on unseen data.
4 GENERALIZATION GAP
While Section 3 established tight stability bounds for robust distributed learning, an open question is how the uncovered stability gap translates into generalization error under the two threat models. To address this, we focus on the smooth convex setting with $\mathrm{SMEA}$ and construct a data distribution where the generalization error is not merely upper bounded by uniform stability but is in fact proportional to it—independently of the threat model.
Lemma 4.1.
Consider the setting described in Section 2, with $m=1$, regardless of the assumed threat model. There exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, data distributions ${\{p_{i}\}}_{i\in\mathcal{H}}$ over $\mathcal{Z}$, such that for any distributed algorithm $\mathcal{A}$, we have:
$\left|\mathbb{E}_{\mathcal{A},\mathcal{S}\sim\otimes_{i\in\mathcal{H}}\left(p_{i}^{\otimes m}\right)}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]\right|\\ =\frac{1}{4(n-f)}\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}\left[\ell(\mathcal{A}(S);z)-\ell(\mathcal{A}(S^{\prime});z)\right],$ (10)
where the dependence on the threat model arises solely through the stability term.
This construction unifies the lower-bound settings (i.e. honest datasets and loss functions) shared by both worst-case data poisoning and Byzantine attack scenarios. As a result, our findings reveal a fundamental separation: there exist data distributions and Byzantine attacks whose generalization error is strictly larger (in terms of convergence rate) than that of the worst-case data poisoning attack.
Theorem 4.2.
Consider the setting of Lemma 4.1, $\frac{n}{3}\leq f<\frac{n}{2}$, $\mathcal{A}=\mathrm{GD}$ with $\mathrm{SMEA}$. There exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, data distributions ${\{p_{i}\}}_{i\in\mathcal{H}}$ over $\mathcal{Z}$ and a Byzantine attack such that for any data poisoning attacks, we have
$\frac{\mathcal{E}^{\mathrm{byz}}_{gen}}{\mathcal{E}^{\mathrm{pois}}_{gen}}\in\Omega\left(\frac{n-f}{\sqrt{f(n-2f)}}\right),$ (11)
where $\mathcal{E}^{\mathrm{byz}}_{gen}$ and $\mathcal{E}^{\mathrm{pois}}_{gen}$ denote generalization errors under the Byzantine and data poisoning attacks, respectively.
In what follows, we complement our theoretical results with numerical experiments that validate the analysis and provide further insights. Then, we discuss the broader intuition underlying the emergence of this generalization gap between the two threat models.
Numerical validation. We numerically instantiate the construction used in proving the stability lower bound under data poisoning, which coincides with the Byzantine case when $f\geq\frac{n}{3}$ (see Sections D.2.1 and C.3 for further details), with $m=1$. Under this configuration, we can use Lemma 4.1 as demonstrated in Theorem 4.2. For our experiments, we fix the learning rate and the Lipschitz coefficient $\gamma=C=1$, the number of epochs $T=5$, the number of total workers $n=15$ and vary the number of misbehaving workers $f$ from $1$ to $7$. As expected, we observe that under data poisoning, our theoretical upper bound (scaled by the factor $\frac{1}{4(n-f)}$ from eq. 10) closely matches the empirically measured generalization error (Figure 1), confirming the practical tightness of our analysis in this setting, including the constant factors.
Next, we evaluate the system’s performance under the Byzantine attack described in the proof of the stability lower bound in Theorem 3.2. In particular, for each $f$, we fix $\mathcal{H}$ to maximize honest gradients’ variance. The Byzantine workers then accelerate the divergence of the parameters, either in the positive or negative direction, triggered by a specific event that depends on the observed communication. Numerically, this is achieved by adaptively crafting updates that remain within the aggregation rule’s selection range (see Section E.2 for details). We observe that this Byzantine attack results in a significantly higher generalization error that grows with the number of Byzantine workers. Importantly, when $f\geq\frac{n}{3}$, the attack produces an error greater than the lower bound reported in Theorem 3.2, scaled by $\frac{1}{4(n-f)}$. This was expected as our data distribution produces datasets matching the worst-case one crafted for the Byzantine lower bound. The gap between the two curves illustrates that our attack optimizes the numerical constant in the lower bound.
Interestingly, even when $f<\frac{n}{3}$, Byzantine failures continue to induce substantially greater generalization error than worst-case data poisoning, emphasizing that the gap persists even with a moderate fraction of misbehaving workers. In fact, as $f<\frac{n}{3}$ decreases, it becomes increasingly challenging (but not impossible) to simultaneously maintain maximal honest gradients’ variance and amplify the influence of Byzantine workers. In the absence of an order-of-magnitude larger lower bound under Byzantine failures in the regime $f<n/3$, it is natural to ask whether the generalization gap also extends to this low-to-moderate fraction of misbehaving workers. While our numerical experiments support this intuition, the open question remains whether the gap is merely a constant factor or truly an order of magnitude.
In summary, our numerical experiment demonstrates that Byzantine failures can have a significantly greater generalization error than worst-case data poisoning across all regimes, from low to high fractions of misbehaving workers. Remarkably, this also confirms the fundamental generalization gap between Byzantine and data poisoning threat models, as highlighted above in Theorem 4.2.
Figure 1: Stability Under Optimal Poisoning And Tailored Byzantine Attacks.
Why threat model differences affect generalization, but not optimization? In contrast to Farhadkhani et al. (2024b), who show similar optimization error guarantees for both threat models, our work reveals a gap in generalization error. The key intuition is as follows. Optimization analysis focuses on how accurately the descent direction is estimated at each step, whereas generalization analysis examines the algorithm’s sensitivity to individual training samples, independent of how informative the estimated direction is for optimization. In the worst case, both corruptions similarly impair optimization direction estimation. However, as discussed in Section 3, stability analysis benefits from the additional regularity and non-adaptive nature of data poisoning. Furthermore, prior work (Allouah et al., 2023a) shows that for the optimization error, $(f,\kappa)$-robustness with $\kappa\in\mathcal{O}(f/n)$ suffices to achieve an upper bound under Byzantine failures that matches the lower bound for poisoned data (Allouah et al., 2023a, Theorem 1 & Proposition 1). Hence, this bound cannot be improved using the regularity of corrupted gradients. We provide further discussion in Appendix H.
5 RELATED WORK
Minimizing empirical risk over honest workers in the presence of Byzantine workers has been extensively studied (Guerraoui et al., 2018; Karimireddy et al., 2021, 2022; Allouah et al., 2023a; Gorbunov et al., 2023; Allouah et al., 2023b), leading to robust aggregation schemes enabling distributed $\mathrm{SGD}$ to attain optimal error even under data heterogeneity. A key contribution of this line of work is the formal analysis of aggregation rules previously studied under i.i.d. assumptions. However, these analyses focus on optimization error and offer limited insight into the generalization performance.
Empirical studies (Baruch et al., 2019; Xie et al., 2020; Shejwalkar and Houmansadr, 2021; Allen-Zhu et al., 2020; Allouah et al., 2023a) have shown that the generalization error under Byzantine attacks is often worse than under data poisoning, but offer no formal explanation for this gap. These findings stand in contrast to theoretical analyses proving comparable statistical error rates across the two threat models (Alistarh et al., 2018; Zhu et al., 2023; Yin et al., 2018). However, these theoretical results assume an i.i.d. data setting, where all honest workers sample from the same distribution, and do not account for data heterogeneity common in distributed learning. We bridge this gap between empirical observations and theoretical results by developing the first theoretical framework to analyze the generalization error of robust first-order iterative methods under heterogeneous data and the two considered threat models. Our findings apply to a much larger class of loss functions, and indeed show that Byzantine attacks can be more harmful to generalization than data poisoning.
A prior work (Farhadkhani et al., 2022b), which claims an equivalence between the two threat models in the context of PAC learning, considers only the classic non-robust distributed $\mathrm{SGD}$ algorithm. Their analysis does not apply to robust distributed $\mathrm{GD}$ or $\mathrm{SGD}$, as classic averaging has no optimization guarantees in presence of misbehaving workers. Another work (Farhadkhani et al., 2024b), which analyzes the relationship between data poisoning and Byzantine failures for robust distributed $\mathrm{SGD}$, considers a streaming setting in which honest workers acquire i.i.d. training samples from their respective data-generating distributions at each iteration. While they show that, in that setting, the two threat models yield comparable population risk (asymptotically), their results and proof techniques do not apply to our more pragmatic setting wherein the training datasets across the honest workers are sampled a priori.
Our work relates to Ye and Ling (2025), which studies uniform stability of robust decentralized $\mathrm{SGD}$ under Byzantine failures. However, their analysis uses a diameter-based robustness definition for aggregation rules. This definition leads to robustness parameters that exhibit suboptimal dependence on the problem dimension or the number of honest agents, as highlighted in (Ye et al., 2024, Table II) and further discussed in (Allouah et al., 2023a, Section 8.3). Consequently, this robustness notion yields suboptimal empirical resilience (cf. Farhadkhani et al. (2022a); Allouah et al. (2023a)), in contrast to covariance-based definitions, such as $(f,\kappa)$-robustness, that we consider. Moreover, Ye and Ling (2025) does not provide any insights into the data poisoning threat model.
6 CONCLUSION
We have rigorously proved, for the first time, a fundamental gap in the generalizability of robust distributed learning under two central threat models: Byzantine failures and data poisoning. Specifically, we have shown that Byzantine failures can cause strictly worse generalization than data poisoning—even when both attacks are optimally designed. Our results explain prior empirical observations overlooked by optimization analyses, revealing that guarantees differ across the two threat models. This insight opens new avenues to better understand robust distributed learning and adapt robustness properties to generalization. In particular, incorporating stability into the design of robust aggregation rules could help achieve not only strong optimization guarantees but also lower population risk.
Practical implications.
Our theoretical results highlight the relevance of cryptographic tools—particularly zero-knowledge proofs (ZKPs) (Goldwasser et al., 1989; Thaler, 2022)—for distributed learning. ZKPs can prevent Byzantine failures by verifying that each worker’s input data lies within an appropriate range and that their model updates remain consistent with their initially committed dataset, all without revealing the data itself (Sabater et al., 2022; Abbaszadeh et al., 2024; Shamsabadi et al., 2024). However, they do not guard against data poisoning, which stems from malicious or corrupted local data. Therefore, cryptographic safeguards must be paired with robust learning algorithms that can tolerate compromised training data to ensure end-to-end resilience.
Future work.
Promising directions include: (i) Identify a robust aggregation rule preserving the loss function’s co-coercivity, or prove this is impossible; (ii) Go beyond worst-case scenarios with data-dependent refinements like those based on on-average stability (Lei and Ying, 2020; Sun et al., 2024; Ye et al., 2025); (iii) Extend results to the locally-poisonous setting (Farhadkhani et al., 2024b); (iv) Investigate whether the trilemma between robustness, privacy and optimization error shown in recent work (Allouah et al., 2023b) also holds for generalization error; (v) Quantify the stability of other first-order algorithms, including momentum and communication-efficient methods (see Appendix H for a discussion).
Acknowledgements
The work of Thomas Boudou and Aurélien Bellet is supported by grant ANR 22-PECY-0002 IPOP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR.
Supplementary Materials: Byzantine Failures Harm the Generalization of Robust Distributed Learning Algorithms More Than Data Poisoning
The supplementary materials are organized as follows.
  • Appendix A contains deferred proofs for Section 2, focusing on the link between uniform stability and generalization bounds under Byzantine failures and data poisoning.
  • Appendix B summarizes the regularity assumptions and includes key expansivity results.
  • Appendix C provides detailed proofs of Theorems 3.1 and 3.2 presented in Section 3.1.
  • Appendix D provides detailed proofs of Theorems 3.3 and 3.4 presented in Section 3.2.
  • Appendix E specifies details supporting Section 4, including generalization error computations (Section E.1) and numerical experiments (Section E.2).
  • Appendix F gives example of results with refined bounded heterogeneity and bounded variance assumptions instead of the more general bounded gradients assumption.
  • Appendix G states high-probability excess risk bounds for robust distributed learning under smooth and strongly convex loss functions.
  • Appendix H discusses additional important points not included in the main text.
Appendix A STABILITY IMPLIES GENERALIZATION UNDER BYZANTINE FAILURES OR DATA POISONING
In this section, we focus on the formal link between uniform stability and generalization under Byzantine failures and data poisoning. The proof follows classic arguments from Bousquet and Elisseeff (2002).
Proposition 2.1.
Consider the setting described in Section 2, under either Byzantine failures or data poisoning. let $\mathcal{A}$ an $\varepsilon$-uniformly stable (randomized) distributed algorithm. Then,
$|\mathbb{E}_{\mathcal{S},\mathcal{A}}[R_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))-\widehat{R}_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))]|\leq\varepsilon.$
Proof.
Recall that $\mathcal{S}$ denotes the collective dataset of honest workers. Let $\mathcal{S}^{\prime}=\{z^{\prime(i,j)},i\in\mathcal{H},j\in\{1,\ldots,m\}\}$ be another independently sampled dataset for honest workers where for all $i\in\mathcal{H}$, $\mathcal{D}^{\prime}_{i}=\{z^{\prime(i,1)},\ldots,z^{\prime(i,m)}\}$ is composed of $m$ i.i.d. data points (or samples) drawn from distribution $p_{i}$. Let $\mathcal{S}^{(i,j)}$ be a neighboring dataset of $\mathcal{S}$ with only a single differing data sample $z^{\prime(i,j)}$ at index $(i,j)\in\mathcal{H}\times\{1,\ldots,m\}$. Then,
$\mathbb{E}_{\mathcal{S},\mathcal{A}}[\widehat{R}_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))]$ $=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime},\mathcal{A}}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\frac{1}{m}\sum_{j=1}^{m}\ell(\mathcal{A}(\mathcal{S}^{(i,j)}),z^{\prime(i,j)})\right]$
$=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime},\mathcal{A}}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\frac{1}{m}\sum_{j=1}^{m}\ell(\mathcal{A}(\mathcal{S}),z^{\prime(i,j)})\right]+\delta$
$=\mathbb{E}_{\mathcal{S},\mathcal{A}}[R_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))]+\delta$
with
$\delta=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime}}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{\mathcal{A}}\left[\ell(\mathcal{A}(\mathcal{S}^{(i,j)}),z^{\prime(i,j)})-\ell(\mathcal{A}(\mathcal{S}),z^{\prime(i,j)})\right]\right]\\ \leq\mathbb{E}_{\mathcal{S}}\mathbb{E}_{\mathcal{S}^{\prime}}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\frac{1}{m}\sum_{j=1}^{m}\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}\left[\ell(\mathcal{A}(\mathcal{S}^{(i,j)}),z)-\ell(\mathcal{A}(\mathcal{S}),z)\right]\right]\leq\varepsilon$
Substituting from the above equation proves the lemma. ∎
Appendix B REGULARITY ASSUMPTIONS AND EXPANSIVITY RESULTS
To analyze the stability of iterative optimization algorithms, we typically rely on regularity assumptions on the loss function. We begin by stating standard regularity assumptions used throughout the paper, along with classical results that follow from them.
Definition B.1 (Regularity assumptions).
Let $\ell:\Theta\to\mathbb{R}$ differentiable, $\Theta\subset\mathbb{R}^{d}$.
  • (i)
    $\ell$ is $C$-Lipschitz-continuous (i.e., $C$-Lipschitz) if there exists $C>0$ such that,
    $\forall u,v\in\Theta,\quad|\ell(u)-\ell(v)|\leq C\|u-v\|_{2}.$
    This property is equivalent to the norm of the gradient of $\ell$ being uniformly bounded by $C$.
  • (ii)
    $\ell$ is $L$-Lipschitz-smooth (i.e., $L$-smooth) if its gradient is $L$-Lipschitz. This is equivalent to the following smoothness inequality:
    $\forall u,v\in\Theta,\quad\ell(u)\leq\ell(v)+\langle\nabla\ell(v),u-v\rangle+\frac{L}{2}\|u-v\|_{2}^{2}.$ (12)
  • (iii)
    $\ell$ is convex if and only if, $\forall u,v\in\Theta$, $\langle\nabla\ell(u)-\nabla\ell(v),u-v\rangle\geq 0.$
    Moreover it is $\mu$-strongly convex if there exists $\mu>0$ such that,
    $\forall u,v\in\Theta,\quad\ell(u)\geq\ell(v)+\langle\nabla\ell(v),u-v\rangle+\frac{\mu}{2}\|u-v\|_{2}^{2}.$
    We note that for a strongly convex function to have bounded gradients, it must be defined on a convex compact set, which then implies boundedness of both the loss and the gradient. To ensure this, we can either penalize the problem or restrict the parameter domain and apply an Euclidean projection at every step. Throughout the paper, we will tacitly assume this when referring to strongly convex functions, which does not limit the applicability of our analysis since the projection does not increase the distance between projected points.
Notably, convex and $L$-smooth functions satisfy an important property known as co-coercivity (see Proposition 5.4 in Bach (2024)). Specifically,
Lemma B.1 (Co-coercivity).
Let $\ell:\Theta\to\mathbb{R}$ differentiable, $\Theta\subset\mathbb{R}^{d}$. If $\ell$ is a convex and $L$-smooth function, then it satisfies the co-coercivity inequality,
$\forall u,v\in\Theta,\quad\langle\nabla\ell(u)-\nabla\ell(v),u-v\rangle\geq\frac{1}{L}\|\nabla\ell(u)-\nabla\ell(v)\|_{2}^{2}.$ (13)
We focus on robust variants of distributed gradient descent and stochastic gradient descent. Below we recall properties of the one-step update. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, $\ell:\Theta\rightarrow\mathbb{R}$ differentiable and $\gamma>0$, we denote:
$G^{\mathcal{A}}_{\gamma}(\theta):=\theta-\gamma\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}g^{(i)}.$
Here, for each $i\in\mathcal{H}$, $g^{(i)}$ denotes the exact gradient $\nabla\ell(\theta)$ in the case of $\mathrm{GD}$, or an unbiased estimate thereof—effectively computed with a sample from the local dataset—in the case of $\mathrm{SGD}$. We recall the expansiveness definition for an update rule from Hardt et al. (2016, Definition 2.3):
Definition B.2 ($\eta$-expansive update rule).
An update rule $G:\Theta\rightarrow\Theta$ is said to be $\eta$-expansive if for all $\theta,\omega\in\Theta$,
$\|G(\theta)-G(\omega)\|_{2}\leq\eta\|\theta-\omega\|_{2}.$
A straightforward extension of the result from Hardt et al. (2016, Lemma 3.6) to the distributed setting yields:
Lemma B.2 (Expansivity of gradient updates).
Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ and $\gamma>0$. Assume $\forall z\in\mathcal{Z},\ell(\cdot,z):\Theta\to\mathbb{R}$ is L-smooth, Then $G^{\mathcal{A}}_{\gamma}$ is $\eta_{G^{\mathcal{A}}_{\gamma}}$-expansive (Hardt et al., 2016, Definition 2.3), with the following expressions:
  • (i)
    $\eta_{G^{\mathcal{A}}_{\gamma}}=(1+\gamma L)$.
  • (ii)
    Assume in addition that $\ell$ is convex. Then, for any $\gamma\leq\frac{2}{L}$, $\eta_{G^{\mathcal{A}}_{\gamma}}=1$.
  • (iii)
    Assume in addition that $\ell$ is $\mu$-strongly convex. Then, for any $\gamma\leq\frac{2}{\mu+L}$, $\eta_{G^{\mathcal{A}}_{\gamma}}=(1-\gamma\frac{L\mu}{\mu+L})$. For $\gamma\leq\frac{1}{L}$, since we have $L\geq\mu$, we can simplify the contractive constant to $\eta_{G^{\mathcal{A}}_{\gamma}}=(1-\gamma\mu)$.
$\eta_{G^{\mathcal{A}}_{\gamma}}$ is referred to as the expansivity coefficient of $\mathcal{A}$ with learning rate $\gamma$.
Appendix C DEFERED PROOFS FROM SECTION 3.1
We first establish supporting results in Section C.1, then prove Theorem 3.1.
C.1 Expectation bounds for honest gradients’ empirical spectral norm
We derive bounds on the expected spectral norm of the empirical covariance matrix of the honest workers’ gradients under the setting of bounded gradient assumption (i.e. Lipschitz-continuity of the loss function).
Lemma C.1.
Consider the setting described in Section 2 under Byzantine failures. Assume the loss function $C$-Lipschitz. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa)$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
$\mathbb{E}_{\mathcal{A}}\|\Sigma_{\mathcal{H},t}\|_{sp}=\mathbb{E}_{\mathcal{A}}[\lambda_{\max}(\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}(g^{(i)}_{t}-\overline{g}_{t}){(g^{(i)}_{t}-\overline{g}_{t})}^{\intercal})]\leq C^{2}.$
Proof.
For the first inequality, we directly use the boundedness of gradients:
$\lambda_{\max}(\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}(g^{(i)}_{t}-\overline{g}_{t}){(g^{(i)}_{t}-\overline{g}_{t})}^{\intercal})=\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\overline{g}_{t}\rangle^{2}\\ \leq\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}\rangle^{2}\leq\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|g^{(i)}_{t}\|^{2}_{2}\leq C^{2}.$
Remark C.1.
In Appendix F, we present example results derived under refined assumptions—namely, bounded heterogeneity and bounded variance—in place of the more general bounded gradients condition (i.e. Lipschitz-continuity of the loss function). These assumptions are more commonly used in the optimization literature (as opposed to the generalization literature), and while they often yield tighter bounds, they do not provide additional conceptual insights within the scope of our discussion.
C.2 Proof of Theorem 3.1
We begin with a unified proof sketch that outlines the key ideas to bound stability under Byzantine failures. We then prove Theorem 3.1.
Proof sketch.
Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ with $F$ a $(f,\kappa)$-robust aggregation rule, and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. Let denote $m^{(i)}_{t}$ and $m^{\prime(i)}_{t}\in\mathbb{R}^{d}$ the local updates provided by worker $i\in\{1,\ldots,n\}$ along the trajectories $t\in\{0,\ldots,T-1\}$. In the context of the uniform algorithmic stability analysis via parameter sensitivity, we track how the parameters diverge along these optimization trajectories. To do so, we decompose the analysis by introducing an intermediate comparison between the robust update $G^{F}_{\gamma_{t}}$ and the averaging over honest workers update $G^{\mathcal{A}}_{\gamma_{t}}$. This enables us to leverage either the regularity of the loss function and the robustness property of the aggregation rule. This approach is primarily motivated by the fact that Byzantine vectors cannot be assumed to exhibit regularities when comparing them.
$\mathbb{E}_{\mathcal{A}}\left[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\right]$ $=\mathbb{E}_{\mathcal{A}}[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})+G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})+G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})-G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}]$
$\leq\underbrace{\mathbb{E}_{\mathcal{A}}\|G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}}_{A}+\underbrace{\gamma_{t}\mathbb{E}_{\mathcal{A}}\left[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})\|_{2}+\|G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\right]}_{B}$
We then bound the term $A$ with classical stability tools from Hardt et al. (2016), and the term $B$ with the $(f,\kappa)$-robust property, the Jensen inequality and the Lemma C.1:
$B\leq 2\gamma_{t}\mathbb{E}_{\mathcal{A}}\sqrt{\kappa\|\Sigma_{\mathcal{H},t}\|_{op}}\leq 2\gamma_{t}\sqrt{\kappa\mathbb{E}_{\mathcal{A}}\|\Sigma_{\mathcal{H},t}\|_{op}}\leq 2\gamma_{t}\sqrt{\kappa}C.$
Theorem 3.1.
Consider the setting described in Section 2 under Byzantine failures. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with $F$ a $(f,\kappa)$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth.
  • (i)
    Assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ convex, with constant learning rate $\gamma\leq\frac{2}{L}$, then we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S^{\prime})};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq 2\gamma C^{2}T\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right).$
  • (ii)
    Assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $\mu$-strongly convex, with constant learning rate $\gamma\leq\frac{1}{L}$, then we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S^{\prime})};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq\frac{2C^{2}}{\mu}\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right).$
Proof.
Let first reason about $\mathcal{A}=\mathrm{GD}$. For each $t\in\{0,\ldots,T-1\}$, following the proof sketch in C.2, we bound the term $\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ where $G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})$. Let $a,b\in\mathcal{H}\times\{1,\ldots,m\}$ denote the indices of the differing samples. For $t\in\{0,\ldots,T-1\}$, we have,
$\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ $=\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}+\frac{\gamma_{t}}{|\mathcal{H}|m}\|\nabla\ell(\theta_{t}^{\prime};z^{\prime(a,b)})-\nabla\ell(\theta_{t}^{\prime};z^{(a,b)})\|_{2}$
$\leq\eta_{G^{\mathrm{GD}}_{\gamma_{t}}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}+\frac{2\gamma_{t}C}{(n-f)m}$
where $\eta^{\mathrm{GD}}_{G{\gamma_{t}}}$ denotes the expansive coefficient, which depends on the regularity properties of the loss function (see Lemma B.2). Let define $\sigma^{F}_{\mathrm{GD}}=2C\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)$, we then resolve the recursion by incorporating the impact of the Byzantine failures as demonstrated in the previous proof sketch.
  • (i)
    If the loss function is Lipschitz, smooth, and convex, then for all $t\in\{0,\ldots,T-1\}$:
    $\|\theta_{t+1}-\theta_{t+1}^{\prime}\|_{2}\leq\|\theta_{t}-\theta_{t}^{\prime}\|_{2}+\gamma\sigma^{F}_{\mathrm{GD}}$
    To conclude the proof, we assume a constant learning rate $\gamma\leq\frac{1}{L}$ and sum the inequality over $t\in{0,\ldots,T-1}$, noting that $\theta_{0}=\theta_{0}^{\prime}$. We then invoke the Lipschitz continuity assumption.
  • (ii)
    If the loss function is Lipschitz, smooth, and strongly convex, then for all $t\in\{0,\ldots,T-1\}$:
    $\|\theta_{t+1}-\theta_{t+1}^{\prime}\|_{2}\leq(1-\gamma\mu)\|\theta_{t}-\theta_{t}^{\prime}\|_{2}+\gamma\sigma^{F}_{\mathrm{GD}}$
    Considering $\gamma\leq\frac{1}{L}$ and summing over $t\in\{0,\ldots,T-1\}$ with $\theta_{0}=\theta_{0}^{\prime}$:
    $\|\theta_{T}-\theta_{T}^{\prime}\|_{2}$ $\leq\gamma\sigma^{F}_{\mathrm{GD}}\sum_{t=0}^{T-1}{\left(1-\gamma\mu\right)}^{t}=\frac{\sigma^{F}_{\mathrm{GD}}}{\mu}\left[1-{\left(1-\gamma\mu\right)}^{T}\right]\leq\frac{\sigma^{F}_{\mathrm{GD}}}{\mu}$
    We conclude the derivation of the bound by invoking the Lipschitz continuity assumption.
For $\mathcal{A}=\mathrm{SGD}$, the same recursion holds for the expected distance $\mathbb{E}_{\mathcal{A}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$ for $t\in\{1,\ldots,T\}$, where $G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}g^{(i)}_{t}$, and for each $i\in\mathcal{H}$, $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i,J^{(i)}_{t})}_{t})$ with $J^{(i)}_{t}$ sampled uniformly from $\{1,\ldots,m\}$. In fact, at each iteration, a single sample is selected uniformly at random from each honest worker. Consequently, with probability $(1-\frac{1}{m})$, the differing samples are not selected,
$\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}=\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2},$
and with probability $\frac{1}{m}$, the differing samples are selected,
$\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ $=\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$
$\qquad+\frac{\gamma_{t}}{|\mathcal{H}|}\|\nabla\ell(\theta_{t}^{\prime};z^{\prime(a,b)})-\nabla\ell(\theta_{t}^{\prime};z^{(a,b)})\|_{2}$
$\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}+\frac{2\gamma_{t}C}{|\mathcal{H}|}.$
We thus obtain the bound:
$\mathbb{E}_{\mathcal{A}}[\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}]\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}[\|\theta_{t}-\theta^{\prime}_{t}\|_{2}]+\frac{2\gamma_{t}C}{(n-f)m}$
Let define $\sigma^{F}_{\mathrm{SGD}}=2C\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)$, we then resolve the recursion, like in the $\mathrm{GD}$ case above, by incorporating the impact of the Byzantine failures as demonstrated in the previous proof sketch:
$\mathbb{E}_{\mathcal{A}}[\|\theta_{t+1}-\theta^{\prime}_{t+1}\|_{2}]\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}[\|\theta_{t}-\theta^{\prime}_{t}\|_{2}]+\gamma_{t}\sigma^{F}_{\mathrm{SGD}}.$ (14)
This conclude the proof. ∎
C.3 Proof of Theorem 3.2
Theorem 3.2.
Consider the setting described in Section 2 under Byzantine failures. Let $\mathcal{A}=\mathrm{GD}$, with $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with the learning rate $\gamma$. Then for both regime stated below, there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
  • (i)
    in the regime of moderate to high fractions of misbehaving workers $\frac{n}{3}\leq f<\frac{n}{2}$,
    $\Omega\left(\gamma C^{2}T\left(\sqrt{\frac{f}{n-2f}}+1+\frac{1}{(n-f)m}\right)\right).$
  • (ii)
    in the regime of low to moderate fractions of misbehaving workers $f<\frac{n}{3}$,
    $\Omega\left(\gamma C^{2}T\frac{f}{n-2f}\right).$
Proof.
The lower bound construction relies on two key elements. First, we maximize the variance among the honest gradients by dividing them into two equal subgroups positioned at the boundary of the Lipschitz constraint. Second, in the regime with a high proportion of Byzantine workers, this setup enables the adversary to completely discard one of the honest subgroups and even outweigh the influence of the other. The extent to which the adversary can bias the aggregate away from the true direction (which is zero by construction) depends on the robustness of the aggregation rule.
In what follows, we provide the proof for the case $d=1$. The result extends naturally to higher dimensions by embedding the one-dimensional construction into $\mathbb{R}^{d}$. We arbitrarily refer to one trajectory as unperturbed, with parameter $\theta_{t}$, and the other as perturbed, with $\theta^{\prime}_{t}$. Let $C,L\in\mathbb{R}^{*}_{+}$, we define the following loss function:
$\text{For }\theta\in\mathbb{R},\text{ }z\in[-C,C],\quad\ell(\theta;z)=z\theta,\quad\ell^{\prime}(\theta;z)=z,\quad\ell^{\prime\prime}(\theta;z)=0.$
By construction, the loss function is convex, $C$-Lipschitz and $L$-smooth. Let $m,n,f\in\mathbb{N}^{*}$ such that $\frac{n}{3}\leq f<\frac{n}{2}$. Let $\gamma$ the learning rate and assume $\theta_{0}=\theta_{0}^{\prime}=0$.
(i) If nfn-f is even and $f\geq\frac{n}{3}$.
We define the following pair of neighboring datasets:
$\begin{split}S=\{D_{1},\ldots,D_{n-f}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n-f\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i\in\mathcal{H}_{+}=\{1,\ldots,\frac{n-f}{2}\},|\mathcal{H}_{+}|=\frac{n-f}{2}\!\!:z^{(i,j)}=C\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in\mathcal{H}_{-}=\{\frac{n-f}{2}+1,\ldots,n-f\},|\mathcal{H}_{-}|=\frac{n-f}{2}\!\!:z^{(i,j)}=-C.\end{split}$
For $t\in\{0,\ldots,T-1\}$, for $i\in\mathcal{H}_{-}$, $g^{(i)}_{t}=g^{(i)\prime}_{t}=-C$, for $i\in\mathcal{H}_{+}\setminus\{1\}$, $g^{(i)}_{t}=g^{(i)\prime}_{t}=C$, $g_{t}^{(1)}=C$ and $g_{t}^{(1)\prime}=\frac{m-1}{m}C<C$. We consider $f$ Byzantine workers, and let denote $\mathcal{B}=\{n-f+1,\ldots,n\}$. Since they can observe all communications, they may adapt their strategy based on the message sent by worker $1$. We define their behavior as follows:
if $g_{1}\leq C$ then
 return an arbitrarily high value (for instance $n\beta$)
else
 return $\beta$ (to be specified below)
Algorithm 1 Byzantine worker behavior
In the first case, the subset selected by the $\mathrm{SMEA}$ aggregation rule consists entirely of honest workers. Consequently, we have $\theta_{T}^{\prime}=\frac{\gamma CT}{m(n-f)}$. In the second case, we choose
$\beta=\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}\right)C,$
such that the $\mathrm{SMEA}$ aggregation rule selects the subset $\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}$, where $\text{Select}_{k}(E)$ denote any subset of size $k$ from a set $E$, i.e., all the Byzantine workers together with n2fn-2f workers from $\mathcal{H}_{+}$. This is made possible because:
[…]
  • $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}})=C^{2}$, and any subset $\text{Select}_{a}(\mathcal{H}_{-})\cup\text{Select}_{b}(\mathcal{H}_{+})\cup\text{Select}_{n-f-a-b}(\mathcal{B})$, with $0<a<\frac{n-f}{2}$, $0\leq b<\frac{n-f}{2}$ yields higher variance.
  • As $f\geq n/3$, $f\geq\frac{n-f}{2}\geq n-2f$. Hence, either $f=\frac{n-f}{2}$ and
    $\mathcal{H}_{+}\cup\text{Select}_{\frac{n-3f}{2}}(\mathcal{B})=\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}=\mathcal{H}_{+}\cup\mathcal{B},$
    or $f>\frac{n-f}{2}$ and $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}_{+}\cup\text{Select}_{\frac{n-3f}{2}}(\mathcal{B})})>\text{Var}({\{g^{(i)}_{t}\}}_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}})$. Hence, the aggregation rule will always pick a larger subgroup over a smaller one (it minimizes the variance magnitude).
  • $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}})=\frac{C^{2}}{4}<C^{2}=\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}})$. In fact,
    $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}})$ $=\frac{f}{n-f}{\left[\beta-\frac{1}{n-f}\left(f\beta+(n-2f)C\right)\right]}^{2}$
    $\qquad+\frac{n-2f}{n-f}{\left[C-\frac{1}{n-f}\left(f\beta+(n-2f)C\right)\right]}^{2}$
    $=\frac{f(n-2f)}{{(n-f)}^{2}}{\left(\beta-C\right)}^{2}$
    $=\frac{f(n-2f)}{{(n-f)}^{2}}C^{2}{\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}-1\right)}^{2}$
    $=\frac{C^{2}}{4}.$
This reasoning applies to all iterations because the gradients, by construction, stay constant during the entire optimization process. As a result,
$\theta_{T}$ $=\theta_{0}-\gamma T\frac{1}{n-f}\sum_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}}g^{(i)}_{0}$
$=-\frac{\gamma T}{n-f}\left(fC\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}\right)+(n-2f)C\right)$
$=-\gamma TC\left(1+\frac{1}{2}\sqrt{\frac{f}{n-2f}}\right).$
Hence,
$\sup_{z\in[-C,C]}|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|$ $=\sup_{z\in[-C,C]}z\left(|\theta_{T}-\theta^{\prime}_{T}|\right)$
$=\gamma C^{2}T\left(1+\frac{1}{2}\sqrt{\frac{f}{n-2f}}+\frac{1}{(n-f)m}\right)\in\Omega\left(\gamma C^{2}T\left(\sqrt{\frac{f}{n-2f}}+\frac{1}{(n-f)m}\right)\right).$
(ii) If nfn-f is odd and $f\geq\frac{n}{3}$.
The same construction works with a slight adaptation. That is, we define the following pair of neighboring datasets:
$\begin{split}S=\{D_{1},\ldots,D_{n-f}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n-f\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i\in\mathcal{H}_{+}=\{1,\ldots,\frac{n-f-1}{2}\},|\mathcal{H}_{+}|=\frac{n-f-1}{2}\!\!:z^{(i,j)}=C\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in\mathcal{H}_{-}=\{\frac{n-f-1}{2}+1,\ldots,n-f-1\},|\mathcal{H}_{-}|=\frac{n-f-1}{2}\!\!:z^{(i,j)}=-C.\\ &\quad z^{(n-f,j)}=0.\end{split}$
Then,
$\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}})=(1-\frac{1}{n-f})C^{2}>\frac{C^{2}}{3}\text{ as }n-f>\frac{n}{2}\geq\frac{3}{2}.$
The remainder of the proof when nfn-f is odd proceeds identically to the previous case when nfn-f is even.
(iii) In the regime $f<n/3$.
For values restricted to $z\in[-C,C]$, variance is maximized by placing all points at the endpoints $\pm C$, with counts as balanced as possible. Thus we take
$k=\left\lceil\frac{n-f}{2}\right\rceil\quad\text{points at }+C,\qquad\left\lfloor\frac{n-f}{2}\right\rfloor=n-f-k\quad\text{points at }-C.$
We have two cases, if nfn-f is even then the variance equals $C^{2}$, else it equals $C^{2}\left(1-\frac{1}{{(n-f)}^{2}}\right)$. Because $f<n/3$ we have n2f>(nf)/2n-2f>(n-f)/2. Hence, we would like to consider the subset $S$ composed by $f$ times $\beta$, $k$ times $C$ and n2fkn-2f-k times C-C. We compute the competing set variance by the identity $\mathrm{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$,
$\mathrm{Var}(S)=\frac{(n-2f)C^{2}+f\beta^{2}}{n-f}-\frac{\big((2k-n+2f)C+f\beta\big)^{2}}{(n-f)^{2}}.$
Requiring $\mathrm{Var}(S)\leq\mathrm{Var}(\mathcal{H})=C^{2}\!\left(1-\frac{(2k-n+f)^{2}}{(n-f)^{2}}\right)$ yields the quadratic inequality in $\beta$
$(n-2f)\beta^{2}+2(n-2f-2k)C\beta-(4k-n+2f)C^{2}\leq 0.$
The discriminant simplifies exactly to $16C^{2}k^{2}$, so the two roots are
$\beta_{-}=-C,\qquad\beta_{+}=\frac{4k-(n-2f)}{n-2f}C=(\frac{4k}{n-2f}-1)C.$
Since (n2f)>0(n-2f)>0, the admissible $\beta$ lie between the two roots. Moreover, since $\frac{2k}{n-2f}=\frac{2\left\lceil\frac{n-f}{2}\right\rceil}{n-2f}\geq\frac{n-f}{n-2f}$, we have $\frac{4k}{n-2f}-1\geq 2\frac{n-f}{n-2f}-1=\frac{2f}{n-2f}+1>1$.
Next, we have:
$\theta_{T}=-\gamma CT\frac{2k-n+2f+\frac{4fk}{n-2f}-f}{n-f}=-\gamma CT\left(\frac{4fk}{(n-f)(n-2f)}+\frac{2k-(n-f)}{n-f}\right),$
with $2k\geq n-f$. Hence, similarly to the previous cases, we have $\theta_{T}^{\prime}>0$, resulting in:
$\sup_{z\in[-C,C]}|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|=\sup_{z\in[-C,C]}z\left(|\theta_{T}-\theta^{\prime}_{T}|\right)\geq\gamma C^{2}T\frac{2f}{n-2f}$
Appendix D DEFERED PROOFS FROM SECTION 3.2
In the following section, we establish uniform algorithmic upper bounds and lower bounds under data poisoning, leveraging the structure of the $\mathrm{SMEA}$ robust aggregation rule.
D.1 Proof of Theorem 3.3
Theorem 3.3.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with $\mathrm{SMEA}$, for $T\in\mathbb{N}^{*}$ iterations. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth.
  • (i)
    If we assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ convex, with constant learning rate $\gamma\leq\frac{2}{L}$, we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S)};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq 2\gamma C^{2}T\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right).$
  • (ii)
    If we assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $\mu$-strongly convex, with constant learning rate $\gamma\leq\frac{1}{L}$, we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S)};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq\frac{2C^{2}}{\mu}\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right).$
Proof.
Let first reason about $\mathcal{A}=\mathrm{GD}$. Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathrm{GD}$ with aggregation rule $\mathrm{SMEA}$ and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. The proof proceeds, by analyzing our robust aggregation as a classical gradient descent, albeit with a random data-dependent subset selection at each iteration. Notably, the poisoned vectors are treated as true gradients, although computed on arbitrary data. For $t\in\{0,\ldots,T-1\}$, let denote,
$S^{*}_{t}\in\underset{S\subseteq\{1,\ldots,n\},|S|=n-f}{\text{argmin}}\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}\left(g^{(i)}_{t}-\overline{g}_{S}\right){\left(g^{(i)}_{t}-\overline{g}_{S}\right)}^{\intercal}\right)$
with $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g^{(i)}_{t}$, where $g^{(i)}_{t}=\nabla\widehat{R}_{i}(\theta_{t})$ and the prime notation $(\cdot)$$^{\prime}$ denotes the corresponding quantities for $\mathcal{S}^{\prime}$. At each step $t\in\{0,\ldots,T-1\}$, the intersection and differences of the two subsets $S^{*}_{t}$ and $S^{*\prime}_{t}$ verifies the following properties:
$n-2f\leq|S^{*}_{t}\cap S^{\prime*}_{t}|\leq n-f\quad\text{and}\quad|S^{*}_{t}\setminus S^{\prime*}_{t}|+|S^{\prime*}_{t}\setminus S^{*}_{t}|\leq 2f,$
and let denote,
$G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})=\theta_{t}-\frac{\gamma_{t}}{n-f}\sum_{i\in S^{*}_{t}\cap S^{\prime*}_{t}}\nabla\widehat{R}_{i}(\theta_{t}),$
and
$G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})=\theta^{\prime}_{t}-\frac{\gamma_{t}}{n-f}\sum_{i\in S^{*}_{t}\cap S^{\prime*}_{t}}\nabla\widehat{R}_{i}(\theta_{t}^{\prime}).$
Let $i\notin\mathcal{H}$, the misbehaving vectors are gradients computed on the true loss function. This enables us to exploit their regularities if $i\in S^{*}_{t}\cap S^{\prime*}_{t}$ or their boundedness if $i\in S^{*}_{t}\setminus S^{\prime*}_{t}\cup S^{\prime*}_{t}\setminus S^{*}_{t}$. This is to be contrasted with the previously studied Byzantine failures—where an intermediate comparison was introduced—as we adopt a more direct bounding technique when subject to data poisoning. Let $a,b\in\mathcal{H}\times\{1,\ldots,m\}$ the indices of the differing samples. Using these relations, we derive the following bound on $\delta_{t+1}=\|\theta_{t+1}-\theta^{\prime}_{t+1}\|_{2}$:
$\delta_{t+1}\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}\nabla\widehat{R}_{i}(\theta_{t})-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}\nabla\widehat{R}_{i}(\theta_{t}^{\prime})\|_{2}$
$\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{(n-f)m}\|\nabla\ell(\theta_{t},z^{(a,b)})-\nabla\ell(\theta_{t}^{\prime},z_{t}^{\prime(a,b)})\|_{2}$
$\quad+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}\nabla\ell(\theta_{t},z_{t}^{(i)})-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}\nabla\ell(\theta_{t}^{\prime},z_{t}^{\prime(i)})\|_{2}$
$\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}C\frac{f+\frac{1}{m}}{n-f}.$
Hence, by Lemma B.2, we obtain the following recursion:
$\delta_{t+1}\leq\eta^{\mathrm{GD}}_{G_{\gamma_{t}}}\delta_{t}+2\gamma_{t}\frac{f+\frac{1}{m}}{n-f}C,$
where we bound
$\eta^{\mathrm{GD}}_{G_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}}=\frac{|S^{*}_{t}\cap S^{\prime*}_{t}|}{n-f}\eta^{\mathrm{GD}}_{G_{\gamma_{t}}}\leq\eta^{\mathrm{GD}}_{G_{\gamma_{t}}},$
since $|S^{*}_{t}\cap S^{\prime*}_{t}|\leq n-f$. The precise value of the expansivity coefficient depends on the regularity assumption on the loss function. We then solve the recursion depending on the regularity of the underlying loss function:
  • (i)
    In the Lipschitz, smooth, and convex case, we have for $t\in\{0,\ldots,T-1\}$:
    $\delta_{t+1}\leq\delta_{t}+2\gamma C\frac{f+\frac{1}{m}}{n-f}$
    We finish the proof by considering a constant learning rate $\gamma\leq\frac{1}{L}$ and by summing over $t\in\{0,\ldots,T-1\}$ with $\theta_{0}=\theta_{0}^{\prime}$. We then invoke the Lipschitz continuity assumption.
  • (ii)
    In the Lipschitz, smooth and strongly convex case, we have for $t\in\{0,\ldots,T-1\}$:
    $\delta_{t+1}\leq(1-\gamma\mu)\delta_{t}+2\gamma C\frac{f+\frac{1}{m}}{n-f}$
    Considering $\gamma\leq\frac{1}{L}$ and summing over $t\in\{0,\ldots,T-1\}$ with $\theta_{0}=\theta_{0}^{\prime}$:
    $\delta_{T}\leq 2\gamma C\frac{f+\frac{1}{m}}{n-f}\sum_{t=0}^{T-1}{\left(1-\gamma\mu\right)}^{t}=2\gamma C\frac{f+\frac{1}{m}}{n-f}\frac{1}{\gamma\mu}\left[1-{\left(1-\gamma\mu\right)}^{T}\right]\leq\frac{2C}{\mu}\frac{f+\frac{1}{m}}{n-f}$
    We conclude the derivation of the bound by invoking the Lipschitz continuity assumption.
For $\mathcal{A}=\mathrm{SGD}$, the same recursion holds for the expected distance $\delta_{t}=\mathbb{E}_{\mathcal{A}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$ for $t\in\{1,\ldots,T\}$. We use similar notation as above, where, for $t\in\{1,\ldots,T\}$, $G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}g^{(i)}_{t}$, for each $i\in\mathcal{H}$, $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i,J^{(i)}_{t})}_{t})$ with $J^{(i)}_{t}$ sampled uniformly from $\{1,\ldots,m\}$, and for each $i\notin\mathcal{H}$, $g^{(i)}_{t}$ is a gradient computed on poisoned data. In fact, at each iteration, a single sample is selected uniformly at random from each worker. Consequently, with probability $1-\frac{1}{m}$:
$\delta_{t+1}$ $\leq\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}g^{(i)}_{t}-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}g^{\prime(i)}_{t}\|_{2}$
$\leq\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}C\frac{f}{n-f},$
and with probability $\frac{1}{m}$:
$\delta_{t+1}\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}g^{(i)}_{t}-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}g^{\prime(i)}_{t}\|_{2}$
$\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\nabla\ell(\theta_{t},z^{(a,b)})-\nabla\ell(\theta_{t}^{\prime},z_{t}^{\prime(a,b)})\|_{2}$
$\quad+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}g^{(i)}_{t}-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}g^{\prime(i)}_{t}\|_{2}$
$\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}C\frac{f+1}{n-f},$
where the second inequality corresponds to the worst-case scenario, namely when the index of the worker with the differing data samples lies in the intersection of the two selected subsets of workers, i.e., $a\in S^{*}_{t}\cap S^{\prime*}_{t}$. Hence, with similar derivation as in the $\mathrm{GD}$ case, we obtain the following recursion:
$\mathbb{E}_{\mathcal{A}}\delta_{t+1}\leq\eta^{\mathrm{SGD}}_{G_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}\delta_{t}+2\gamma_{t}\frac{f+\frac{1}{m}}{n-f}C.$ (15)
We finish the proof as above, the depending on the regularity of the underlying loss function. ∎
Remark D.1.
Our bounds does not depend on the robustness coefficient $\kappa$ of the $\mathrm{SMEA}$ robust aggregation, but solely on its structural characteristic—namely, the sub-sampling of worker vectors. Additionally, our proof does not exploit the randomness in the selection of workers at each step, as this would require assumptions regarding the local distributions of the misbehaving workers to evaluate the probability of selecting the differing data samples’ gradients. Interestingly, in practice, one could observe improved stability compared to the setting without misbehaving workers, possibly due to an extended “burn-in period”. That is, if the differing samples are not selected by the aggregation rule when first drawn. This phenomenon could be further explored through numerical experiments.
Remark D.2.
We can actually achieve the same theoretical bound as the non-misbehaving workers framework without the robust aggregation rule. Specifically, by reverting to the classical non-robust mean aggregation, we are able to handle terms as described in Hardt et al. (2016), since data heterogeneity does not manifest in the worst-case uniform algorithmic stability analysis. However, it is important to note that while this approach simplifies the analysis, the population risk may still remain unbounded due to the inherently unbounded optimization error that can arise from mean aggregation when facing an arbitrarily poisoned dataset.
D.2 Proof of Theorem 3.4 for GD
The following lemma plays a key role in understanding the $\mathrm{SMEA}$ aggregation rule, which is essential for the subsequent lower-bound constructions.
Lemma D.1.
Let $n,f\in\mathbb{N}^{*},f<\frac{n}{2}$ and $g_{A},g_{B},g_{C}$ three collinear vectors in $\mathbb{R}^{d},d\geq 1$. Assume we observe a pool of $n$ vectors, consisting of $\alpha$ copies of $g_{A}$, $\beta$ copies of $g_{B}$ and $\gamma$ copies of $g_{C}$, where $\alpha,\beta,\gamma\in\{0,\ldots,n\}$, $\alpha+\beta+\gamma=n$. We abstract each subset $S\subseteq\{1,\ldots,n\}$, composed of |S|=nf|S|=n-f vectors from our pool of vectors as $S_{a,b,c}$ with $a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},c\in\{0,\ldots,\gamma\}$, such that a+b+c=nfa+b+c=n-f and $S$ is composed by $a$ copies of $g_{A}$, $b$ copies of $g_{B}$ and c copies of $g_{C}$. With this notation, we have:
$S^{*}$ $\in\underset{\begin{subarray}{c}S\subset\{1,\ldots,n\},\\ |S|=n-f\end{subarray}}{\text{argmin}}\lambda_{\max}(\frac{1}{|S|}\sum_{i\in S}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal})$
$=\frac{1}{{(n-f)}^{2}}\underset{\begin{subarray}{c}S_{a,b,c}\\ a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},\\ c\in\{0,\ldots,\gamma\},\text{ }a+b+c=n-f\end{subarray}}{\text{argmin}}\{ab\|g_{A}-g_{B}\|_{2}^{2}+bc\|g_{B}-g_{C}\|_{2}^{2}+ac\|g_{A}-g_{C}\|_{2}^{2}\}$
Let $u$ be a unit vector indicating the line the vectors are colinear to. Since all computations are effectively confined to this one-dimensional subspace, we could equivalently emphasize that:
$\lambda_{\max}(\frac{1}{|S|}\sum_{i\in S}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal})=\frac{1}{|S|}\sum_{i\in S}{(g_{i}^{\intercal}u-\overline{g}_{S}^{\intercal}u)}^{2}\lambda_{\max}(uu^{\intercal})=\text{Var}\left({\{g_{i}^{\intercal}u\}}_{i\in S}\right)$
Proof.
Let $S_{a,b,c}$ with $a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},c\in\{0,\ldots,\gamma\}$, such that a+b+c=nfa+b+c=n-f and $S$ is composed by $a$ copies of $g_{A}$, $b$ copies of $g_{B}$ and c copies of $g_{C}$. Then $\overline{g}_{S}=\frac{1}{n-f}(ag_{A}+bg_{B}+cg_{C})$, $g_{A}-\overline{g}_{S}=\frac{b}{n-f}(g_{A}-g_{B})+\frac{c}{n-f}(g_{A}-g_{C})$, $g_{B}-\overline{g}_{S}=\frac{a}{n-f}(g_{B}-g_{A})+\frac{c}{n-f}(g_{B}-g_{C})$ and $g_{C}-\overline{g}_{S}=\frac{a}{n-f}(g_{C}-g_{A})+\frac{b}{n-f}(g_{C}-g_{B})$. We already note that the outer product of vectors $g_{A}$, $g_{B}$ and $g_{C}$ commute because they are collinear. We denote $(*)={(n-f)}^{2}\sum_{i\in S_{a,b,c}}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal}$, by expending and re-arranging the terms:
$(*)$ $=a\times[b^{2}(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}+c^{2}(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}+2bc(g_{A}-g_{B}){(g_{A}-g_{C})}^{\intercal}]$
$\quad+b\times[a^{2}(g_{B}-g_{A}){(g_{B}-g_{A})}^{\intercal}+c^{2}(g_{B}-g_{C}){(g_{B}-g_{C})}^{\intercal}+2ac(g_{B}-g_{A}){(g_{B}-g_{C})}^{\intercal}]$
$\quad+c\times[a^{2}(g_{C}-g_{A}){(g_{C}-g_{A})}^{\intercal}+b^{2}(g_{C}-g_{B}){(g_{C}-g_{B})}^{\intercal}+2ab(g_{C}-g_{A}){(g_{C}-g_{B})}^{\intercal}]$
$=(n-f)ab(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}-abc(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}$
$\quad+(n-f)ac(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}-abc(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}$
$\quad+(n-f)bc(g_{B}-g_{C}){(g_{B}-g_{C})}^{\intercal}-abc(g_{B}-g_{C}){(g_{B}-g_{C})}^{\intercal}$
$\quad+2abc\left[(g_{A}-g_{B}){(g_{A}-g_{C})}^{\intercal}+(g_{B}-g_{A}){(g_{B}-g_{C})}^{\intercal}+(g_{C}-g_{A}){(g_{C}-g_{B})}^{\intercal}\right]$
We then re-arrange the terms and simplify the result:
$(*)$ $=(n-f)ab(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}-abc(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}$
$\quad+(n-f)ac(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}-abc(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}$
$\quad+(n-f)bc(g_{B}-g_{C}){(g_{B}-g_{C})}^{\intercal}-abc(g_{B}-g_{C}){(g_{B}-g_{C})}^{\intercal}$
$\quad+abc\left[(g_{A}-g_{B}){(g_{A}-g_{C})}^{\intercal}-(g_{A}-g_{B}){(g_{B}-g_{C})}^{\intercal}\right]$
$\quad+abc\left[(g_{A}-g_{B}){(g_{A}-g_{C})}^{\intercal}-(g_{A}-g_{C}){(g_{C}-g_{B})}^{\intercal}\right]$
$\quad+abc\left[(g_{B}-g_{A}){(g_{B}-g_{C})}^{\intercal}-(g_{C}-g_{A}){(g_{B}-g_{C})}^{\intercal}\right]$
$=(n-f)\left[ab(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}+ac(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}+bc(g_{B}-g_{C}){(g_{B}-g_{C})}^{\intercal}\right]$
To conclude the proof, we leverage the fact that $g_{A},g_{B},g_{C}$ are collinear vectors. Let $u$ be a unit vector representing the line the vectors are colinear to. Since each $g\in\{g_{A},g_{B},g_{C}\}$ lies along the line spanned by $u$, we can factor out $u$ and use the fact that $\lambda_{\max}(uu^{\intercal})=1$:
$\lambda_{\max}(\frac{1}{n-f}\sum_{i\in S_{a,b,c}}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal})=\frac{1}{{(n-f)}^{2}}\bigg(ab\|g_{A}-g_{B}\|_{2}^{2}+bc\|g_{B}-g_{C}\|_{2}^{2}+ac\|g_{A}-g_{C}\|_{2}^{2}\bigg)$
D.2.1 Convex case
Theorem D.2.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}=\mathrm{GD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{2}{L}$. Then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
$\Omega\left(\gamma C^{2}T\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right)\right).$
Proof.
In what follows, we provide the proof for the case $d=1$. The result extends naturally to higher dimensions by embedding the one-dimensional construction into $\mathbb{R}^{d}$. An explicit construction for $d=2$ is presented in the lower-bound proof for $\mathrm{SGD}$, with $\mathrm{SMEA}$, and convex loss function. We arbitrarily refer to one trajectory as unperturbed, with parameter $\theta_{t}$, and the other as perturbed, with $\theta^{\prime}_{t}$. The proof is structured in the following three parts:
  • (i)
    First, we specify the loss function and the pair of neighboring datasets used. We then describe the structure of our problem, along with its regularity. Broadly speaking, the setup involves three distinct subgroups of workers. The first subgroup acts as a neutral or pivot subgroup that contains the differing samples, while the other two subgroups compete to be selected by the aggregation rule alongside the neutral subgroup.
  • (ii)
    We then show that, within our setup, it is possible to favor one subgroup through initialization and the other through perturbation introduced by the differing sample. This property holds inductively throughout the optimization process, ensuring that the unperturbed and perturbed parameters converge to distinct minimizers. As a result, the two trajectories remain divergent at each iteration, introducing an additional expansive term relative to the behavior observed under standard averaging rule.
  • (iii)
    Finally, we leverage this construction to derive a lower bound on uniform stability, which matches the upper bound established in Theorem 3.3.
Let $C,L\in\mathbb{R}^{*}_{+}$, we define the following loss function:
$\text{For }\theta\in\mathbb{R},\text{ }z\in[-C,C],\quad\ell(\theta;z)=z\theta,\quad\ell^{\prime}(\theta;z)=z,\quad\ell^{\prime\prime}(\theta;z)=0.$
By construction, the loss function is convex, $C$-Lipschitz and $L$-smooth. Let $m,n,f\in\mathbb{N}^{*}$ such that $f<\frac{n}{2}$. We define:
$0<\psi:=\sqrt{\frac{n-2f-\frac{2}{m}}{n-2f+\frac{2}{m}}}<1,$
and the following pair of neighboring datasets:
$\begin{split}S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:z^{(1,j)}=0,\quad z^{(1,1)}=-C,\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:z^{(i,j)}=0.\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:z^{(i,j)}=\frac{1+\psi}{2}C.\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:z^{(i,j)}=-C.\end{split}$
Building on this setup, we first characterize the form of the parameters and gradients throughout the training process. We then specify the conditions under which our lower bound holds. Let $\gamma\leq\frac{2}{L}$ the learning rate and assume $\theta_{0}=\theta_{0}^{\prime}=0$. For $t\in\{0,\ldots,T-1\}$, $g^{(N)}_{t}=g^{(N)\prime}_{t}=g^{(1)\prime}_{t}=0$, $g^{(F)}_{t}=g^{(F)\prime}_{t}=-C$, $g^{(E)}_{t}=g^{(E)\prime}_{t}=\frac{1+\psi}{2}C$, and $g^{(1)}_{t}=-\frac{C}{m}$.
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size nfn-f, and for all $t\in\{0,\ldots,T-1\}$,
$\lambda_{\max}^{S}\vcentcolon=\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}(g^{(i)}_{t}-\overline{g}_{S}){(g^{(i)}_{t}-\overline{g}_{S})}^{\intercal}\right)=\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}(g^{(i)}_{0}-\overline{g}_{S}){(g^{(i)}_{0}-\overline{g}_{S})}^{\intercal}\right)$
with $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g^{(i)}_{0}$. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule:
$S^{*}_{t}\in\underset{S\subseteq\{1,\ldots,n\},|S|=n-f}{\text{argmin}}\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}\left(g^{(i)}_{t}-\overline{g}_{S}\right){\left(g^{(i)}_{t}-\overline{g}_{S}\right)}^{\intercal}\right).$
We highlight that the gradients are independent of the parameters and remain constant throughout the optimization process, as they are equal to the mean of their local samples.
With a slight abuse of notation, we denote, for instance, the set $\{1\}\cup N\cup E$ by $\{1,N,E\}$. To uncover the source of instability, we consider the following two key conditions:
  • (C1)
    By Lemma D.1, for all $t\in\{0,\ldots,T-1\}$, $S^{*\prime}_{t}=\{1,N,E\}$ is equivalent to
    $\lambda_{\max}^{\prime\{1,N,F\}}=\frac{f(n-2f)}{{(n-f)}^{2}}C^{2}>\frac{f(n-2f)}{{(n-f)}^{2}}{\left(\frac{1+\psi}{2}\right)}^{2}C^{2}=\lambda_{\max}^{\prime\{1,N,E\}},$
    which is ensured by construction. In fact, we again recall that the gradients are independent of the parameters and remain constant throughout the optimization process. Next, we demonstrate that any subset composed of a mixture of gradients from 1,N{1,N}, $F$, and $E$ necessarily yields a larger maximum eigenvalue. This is due to the fact that the vectors $g^{(F)\prime}_{0}$ and $g^{(E)\prime}_{0}$ are collinear but oriented in opposite directions, thereby contributing additively to the overall magnitude. To make this precise, we consider the following two cases:
    • (1)
      Let denote $S$ the set were we remove $q<f$ gradients of the subgroup $E$ and add $q$ gradients of the subgroup $F$ from the set $\{1,N,E\}$, we end up comparing the previous quantity to the following, by Lemma D.1:
      $\lambda_{\max}^{\prime S}$ $=\frac{(f-q)(n-2f)}{{(n-f)}^{2}}C^{2}\|g^{(E)\prime}_{0}-g^{(N)\prime}_{0}\|_{2}^{2}+\frac{q(n-2f)}{{(n-f)}^{2}}\|g^{(F)\prime}_{0}-g^{(N)\prime}_{0}\|_{2}^{2}$
      $\quad+\frac{q(f-q)}{{(n-f)}^{2}}\|g^{(F)\prime}_{0}-g^{(E)\prime}_{0}\|_{2}^{2}$
      $=\frac{(f-q)(n-2f)}{{(n-f)}^{2}}C^{2}{\left(\frac{1+\psi}{2}\right)}^{2}+\frac{q(n-2f)}{{(n-f)}^{2}}C^{2}+q(f-q)C^{2}{\left(\frac{3+\psi}{2}\right)}^{2}$
      $=\lambda_{\max}^{\prime\{1,N,E\}}+\frac{q(n-2f)}{{(n-f)}^{2}}C^{2}\left(1-{\left(\frac{1+\psi}{2}\right)}^{2}\right)+\frac{q(f-q)}{{(n-f)}^{2}}C^{2}{\left(\frac{3+\psi}{2}\right)}^{2}$
      $>\lambda_{\max}^{\prime\{1,N,E\}}.$
    • (2)
      Similarly, let denote $S$ the set were we remove $q<\min(f,n-2f)$ gradients of the subgroup $\{1,N\}$ and add $q$ gradients of the subgroup $F$ from the set $\{1,N,E\}$, we end up comparing the previous quantity to the following, by Lemma D.1:
      $\lambda_{\max}^{\prime S}$ $=\frac{f(n-2f-q)}{{(n-f)}^{2}}C^{2}\|g^{(E)\prime}_{0}-g^{(N)\prime}_{0}\|_{2}^{2}+\frac{q(n-2f-q)}{{(n-f)}^{2}}\|g^{(F)\prime}_{0}-g^{(N)\prime}_{0}\|_{2}^{2}$
      $\quad+\frac{qf}{{(n-f)}^{2}}\|g^{(F)\prime}_{0}-g^{(E)\prime}_{0}\|_{2}^{2}$
      $=\lambda_{\max}^{\prime\{1,N,E\}}+\frac{q(n-2f-q)}{{(n-f)}^{2}}C^{2}+\frac{qf}{{(n-f)}^{2}}C^{2}\left({\left(\frac{3+\psi}{2}\right)}^{2}-{\left(\frac{1+\psi}{2}\right)}^{2}\right)$
      $>\lambda_{\max}^{\prime\{1,N,E\}}.$
  • (C2)
    As shown in $(C1)$, the core tension in the choice of the $\mathrm{SMEA}$ aggregation rule ultimately reduces to a comparison between the sets $\{1,N,E\}$ and $\{1,N,F\}$, a consideration that also applies for $(C2)$. By Lemma D.1, for all $t\in\{0,\ldots,T-1\}$,
[…]
    $\left(n-2f-\frac{2}{m}\right)|g^{(F)}_{t}|^{2}<(n-2f+\frac{2}{m})|g^{(E)}_{t}|^{2},$
    which holds by construction as:
[…]
Hence, by construction $\mathrm{SMEA}$ always selects $\{1,N,F\}$ for the unperturbed run, and $\{1,N,E\}$ for the perturbed one. Let denote $p=\frac{f+1/m}{n-f}$, as $\theta_{0}=\theta^{\prime}_{0}=0$, we then have by the $(C1)$ and $(C2)$ conditions,
$\theta^{\prime}_{T}=-\gamma\sum_{t=0}^{T-1}\frac{1}{n-f}\sum_{i\in\{1,N,E\}}g^{(i)\prime}_{t}=-\frac{f}{n-f}\gamma\frac{1+\psi}{2}\gamma CT,$
and
$\theta_{T}=-\gamma\sum_{t=0}^{T-1}\frac{1}{n-f}\sum_{i\in\{1,N,F\}}g^{(i)}_{t}=p\gamma CT.$
Since $\ell(\theta^{\prime}_{T};z)\leq 0$, this directly yields the following lower bound on uniform stability:
$\sup_{z\in[-C,C]}|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|\geq\ell(\theta_{T};C)=C\theta_{T}=p\gamma C^{2}T.\qed$
D.2.2 Strongly convex case
The proof structure with strongly convex objectives closely parallels that of Theorem D.2. We refer to it where appropriate to avoid unnecessary repetition, while explicitly highlighting the key differences.
Theorem D.3.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}=\mathrm{GD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations, $T\geq\frac{\ln\left(1-c\right)}{\ln\left(1-\gamma\mu\right)}$, $0<c<1$, with $\gamma\leq\frac{1}{L}$. Then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and $\mu$-stongly convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
$\Omega\left(\frac{C^{2}}{\mu}\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right)\right).$
Proof.
In what follows, we provide the proof for the case $d=1$. The result extends naturally to higher dimensions by embedding the one-dimensional construction into $\mathbb{R}^{d}$. We arbitrarily refer to one trajectory as unperturbed, with parameter $\theta_{t}$, and the other as perturbed, with $\theta^{\prime}_{t}$. We adopt a reasoning analogous to that used in the lower bound proof in Theorem D.2. As explained in Theorem D.2, we structure our proof into three paragraphs (i), (ii) and (iii) below.
Let $C,L,\mu\in\mathbb{R}^{*}_{+}$, with $\mu\leq L$, we define the following loss function:
$\forall\theta\in\Theta=B(0,\frac{C}{2\mu}),z\in\mathcal{Z}=B(0,\frac{C}{2\mu}),$
$\ell(\theta;z)=\frac{\mu}{2}{(\theta-z)}^{2},\quad\ell^{\prime}(\theta;z)=\mu(\theta-z),\quad\ell^{\prime\prime}(\theta;z)=\mu$
By construction, the loss function is $\mu$-convex, $C$-Lipschitz and $L$-smooth. Let $m,n,f\in\mathbb{N}^{*}$ such that $f<\frac{n}{2}$. We define,
$\max\{\sqrt{\frac{n-2(f+\frac{1}{m})}{n-2(f-\frac{1}{m})}},1-\frac{4}{m(n-2f)}\}<\psi<1,$
and the following pair of neighboring datasets:
$\begin{split}S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{(x^{(1,1)},y^{(1,1)})\}\cup\{(x^{\prime(1,1)},y^{\prime(1,1)})\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{(x^{(i,1)},y^{(i,1)}),\ldots,(x^{(i,m)},y^{(i,m)})\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:z^{(i,j)}=0,\quad z^{(1,1)}=\frac{C}{2\mu},\quad z^{\prime(1,1)}=0.\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:z^{(i,j)}=0.\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:z^{(i,j)}=-\frac{1+\psi}{2}\frac{C}{2\mu}.\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:z^{(i,j)}=\frac{C}{2\mu}.\end{split}$
Let $\gamma\leq\frac{1}{L}$ the learning rate. Following this setting, we first state the form that take the parameters and the gradients during the training process, then the conditions under which we claim our instability: for $t\in\{0,\ldots,T-1\},\theta_{0}=\theta_{0}^{\prime}=0$, $g^{(N)}_{t}=\mu\theta^{\prime}_{t}$ $g^{\prime(N)}_{t}=g^{\prime(1)}_{t}=\mu\theta^{\prime}_{t}$, $g^{(1)}_{t}=(1-\frac{1}{m})g^{(N)}_{t}+g^{(F)}_{t}/m$, $g^{(E)}_{t}=\mu\left(\theta_{t}+\frac{1+\psi}{2}\frac{C}{2\mu}\right)$, $g^{(E)\prime}_{t}=\mu\left(\theta^{\prime}_{t}+\frac{1+\psi}{2}\frac{C}{2\mu}\right)$, $g^{(F)}_{t}=\mu\left(\theta_{t}-\frac{C}{2\mu}\right)$ and $g^{(F)\prime}_{t}=\mu\left(\theta^{\prime}_{t}-\frac{C}{2\mu}\right)$.
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size nfn-f. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule, for all $t\in\{0,\ldots,T-1\}$:
$S^{*}_{t}\in\underset{S\subseteq\{1,\ldots,n\},|S|=n-f}{\text{argmin}}\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}\left(g^{(i)}_{t}-\overline{g}_{S}\right){\left(g^{(i)}_{t}-\overline{g}_{S}\right)}^{\intercal}\right).$
With a slight abuse of notation, we denote, for instance, the set $\{1\}\cup N\cup E$ by $\{1,N,E\}$. To uncover the source of instability, we consider the following two key conditions:
  • (C1)
    As shown in the $(C1)$ condition of Theorem D.2, the core tension in the choice of the $\mathrm{SMEA}$ aggregation rule ultimately reduces to a comparison between the sets $\{1,N,E\}$ and $\{1,N,F\}$, a consideration that also applies here. By Lemma D.1,
    $S^{*\prime}_{0}=\{1,N,E\}\underset{\mathrm{SMEA}}{\Longleftrightarrow}|g^{(F)\prime}_{0}|_{2}=\mu\frac{C}{2\mu}>\mu\frac{1+\psi}{2}\frac{C}{2\mu}=|g^{(E)\prime}_{0}|_{2}.$
    Let denote
    $\lambda_{\max,t}^{\{1,N,E\}\prime}\vcentcolon=\lambda_{\max}\left(\frac{1}{n-f}\sum_{i\in\{1,N,E\}}(g^{(i)\prime}_{t}-\overline{g}^{\prime}_{S}){(g^{(i)\prime}_{t}-\overline{g}^{\prime}_{S})}^{\intercal}\right),$
    and
    $\lambda_{\max,t}^{\{1,N,F\}\prime}\vcentcolon=\lambda_{\max}\left(\frac{1}{n-f}\sum_{i\in\{1,N,F\}}(g^{(i)\prime}_{t}-\overline{g}^{\prime}_{S}){(g^{(i)\prime}_{t}-\overline{g}^{\prime}_{S})}^{\intercal}\right).$
    We then prove that for all $t\geq 0$, $S^{*\prime}_{t}=\{1,N,E\}$, and the unperturbed run converges to the minimizer of the subgroup $E$. For this property to be true, we require:
    $\lambda_{\max,t}^{\{1,N,E\}\prime}<\lambda_{\max,t}^{\{1,N,F\}\prime}.$
    Again, by Lemma D.1, it is equivalent to:
    ${\left(\frac{1+\psi}{2}\right)}^{2}\frac{C^{2}}{4}<\frac{C^{2}}{4},$
    which holds by construction. As the condition is independent of $t$, hence true for all $t$, this conclude the first condition.
  • (C2)
[…]
    $\lambda_{\max,t}^{\{1,N,F\}}<\lambda_{\max,t}^{\{1,N,E\}}.$
    Again, by Lemma D.1, it is equivalent to:
    $f(n-2f-1)\frac{C^{2}}{4}+f\frac{C^{2}}{4}{(1-\frac{1}{m})}^{2}+(n-2f-1)\frac{C^{2}}{4m^{2}}\\ <f(n-2f-1){\left(\frac{1+\psi}{2}\right)}^{2}\frac{C^{2}}{4}+f\frac{C^{2}}{4}{\left(\frac{1+\psi}{2}+\frac{1}{m}\right)}^{2}+(n-2f-1)\frac{C^{2}}{4m^{2}}$
    That is,
    $0<-(n-2f-1)\left(\frac{1-\psi}{2}\right)\left(\frac{3+\psi}{2}\right)+\left(\frac{3+\psi}{2}\right)\left(\frac{2}{m}-\frac{1-\psi}{2}\right),$
    which is equivalent to:
    $\psi>\frac{4}{(n-f)m}.$
    As the condition is independent of $t$, hence true for all $t$, this conclude the second condition.
Let denote $p=\frac{f+1/m}{n-f}$. We have proven that we have for $t\geq 0$:
$\mathrm{SMEA}(g^{(1)}_{t},\ldots,g^{(n)}_{t})=\frac{1}{n-f}\sum_{i\in\{1,N,F\}}g^{(i)}_{t}=\mu\theta_{t}-p\frac{C}{2}$
$\mathrm{SMEA}(g^{(1)\prime}_{t},\ldots,g^{(n)\prime}_{t})=\frac{1}{n-f}\sum_{i\in\{1,N,E\}}g^{(i)\prime}_{t}=\mu\theta^{\prime}_{t}+\frac{f}{n-f}\frac{1+\psi}{2}\frac{C}{2}$
For $t\geq 0$, $\theta_{t+1}=\left(1-\gamma\mu\right)\theta_{t}+\frac{\gamma pC}{2}$ and $\theta^{\prime}_{t+1}=\left(1-\gamma\right)\theta^{\prime}_{t}-\frac{\gamma f(1+\psi)C}{4(n-f)}$. Hence: $\theta^{\prime}_{T}\leq 0$ and,
$\theta_{T}=\frac{pC}{2\mu}\left(1-{\left(1-\gamma\mu\right)}^{T}\right)$
Provided that there exist $0<c<1$ such that $T\geq\frac{\ln\left(1-c\right)}{\ln\left(1-\gamma\mu\right)}$, this implies: $|\theta_{T}-\theta^{\prime}_{T}|\geq|\theta_{T}|\in\Omega(\frac{pC}{\mu})$. Finally, let define $\chi=-\text{sign}(\theta_{T}+\theta^{\prime}_{T})$, we have:
$|\ell(\theta_{T};\chi\frac{C}{2\mu})-\ell(\theta^{\prime}_{T};\chi\frac{C}{2\mu})|$ $=\frac{\mu}{2}|{\left(\theta_{T}-\chi\frac{C}{2\mu}\right)}^{2}-{\left(\theta^{\prime}_{T}-\chi\frac{C}{2\mu}\right)}^{2}|$
$=\frac{\mu}{2}|\theta_{T}-\theta^{\prime}_{T}||\theta_{T}+\theta^{\prime}_{T}+\text{sign}(\theta_{T}+\theta^{\prime}_{T})\frac{C}{\mu}|$
$\geq\frac{\mu}{2}|\theta_{T}|\frac{C}{\mu}\in\Omega(\frac{pC^{2}}{\mu})\qed$
D.3 Proof of Theorem 3.4 for SGD
In this section, we derive a lower bound on the uniform stability of distributed $\mathrm{projected-GD}$ with $\mathrm{SMEA}$ under data poisoning. The proof structure closely parallels that of Theorem D.2. We refer to it where appropriate to avoid unnecessary repetition, while explicitly highlighting the key differences.
Theorem D.4.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}=\mathrm{projected-SGD}$ with $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$. Assume there exist a constant $\tau\geq c>0$ for an arbitrary c, such that $T\geq\tau m$, then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
$\Omega\left(\gamma C^{2}T\left(\frac{f}{n-f}+\frac{1}{(n-f)m}\right)\right).$
Proof.
We adopt a reasoning analogous to that used in the lower bound proof in Theorem D.2. However, we can no longer rely on a linear loss function. The reason is that, in the linear case, the gradients are independent of the model parameters, so when the differing samples are not selected, the perturbed and unperturbed executions remain indistinguishable—thus precluding any divergence. We also have to account for the first occurrence of the differing samples, denoted as $T_{0}=\inf\{t;J^{(1)}_{t-1}=1\}$. In fact, we require to ensure the conditions $(C1)$ (that is $S^{*\prime}_{0}=\{1,N,E\}$) at initialization but we now would like to ensure $(C2)$ when the differing samples are drawn (that is $S^{*}_{t}=\{1,N,F\}$ for $t\geq t_{0}-1$, knowing $T_{0}=t_{0}$ for any time $t_{0}<T$). To do so, we modify the possible parameters we are searching through our optimization process, making the condition at initialization sufficient to ensure $(C1)$ and $(C2)$ for any $t_{0}\in\{0,\ldots,T-1\}$. By defining $\Theta=\{\lambda v;0\leq\lambda<\infty\}$ and applying the projection $p_{\Theta}$ at each step to constrain the parameter within the set of admissible values, we ensure, under condition $(C1)$ that $\theta^{\prime}_{t}=0$ for all $t>0$ and $\theta_{t}=0$ for $t<t_{0}$. Consequently, the condition $(C2)$ is ensured by construction for any time the differing sample is drawn. While using a projection simplifies the lower bound proof (Section D.3), it does not affect the tightness of our result since the upper bound on uniform stability remains valid for $\mathrm{projected}\text{--}\mathrm{SGD}$ as the projection does not increase the distance between projected points.
As explained in Theorem D.2, we structure our proof into three paragraphs (i), (ii) and (iii) below.
Let $C,L\in\mathbb{R}^{*}_{+}$, we define the following dataset and loss:
$\forall\theta\in\Theta=\{\lambda v;\lambda\in\mathbb{R}\}\subset\mathbb{R}^{2},(x,y)\in B(0,\sqrt{L})\times\mathbb{R},$
$\ell(\theta;(x,y))=$ $\begin{cases}\frac{1}{2}{\left(\theta^{T}x-y\right)}^{2}&\text{if }|\theta^{T}x-y|\|x\|_{2}\leq C\\ C\left(\frac{|\theta^{T}x-y|}{\|x\|_{2}}-\frac{C}{2\|x\|^{2}_{2}}\right)&\text{otherwise}.\end{cases}$
$\nabla\ell(\theta;(x,y))=$ $\begin{cases}(\theta^{T}x-y)x&\text{if }|\theta^{T}x-y|\|x\|_{2}\leq C\\ C\text{ sign}(\theta^{T}x-y)\frac{x}{\|x\|_{2}}&\text{otherwise}.\end{cases}$
$\nabla^{2}\ell(\theta;(x,y))=$ $\begin{cases}0\preceq xx^{T}\preceq LI_{d}&\text{if }|\theta^{T}x-y|\|x\|_{2}\leq C\\ 0&\text{otherwise}.\end{cases}$
By construction, the loss function is convex, $C$-Lipschitz and $L$-smooth. Let $m,n,f\in\mathbb{N}^{*}$ such that $f<\frac{n}{2}$. We define,
$0<\psi=\frac{n-2f-2}{n-2f}<1,$
$\min\{1-\psi,\gamma L\frac{f}{n-f}\}>\epsilon>0$
and the following pair of neighboring datasets:
$\begin{split}&\quad v\in B(0,\sqrt{L}),\quad\|v\|^{2}_{2}=L,\quad\beta=\frac{C}{\sqrt{L}},\quad\alpha=(1-\epsilon)\beta,\quad b=\frac{1}{\sqrt{T}}\\ S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{(x^{(1,1)},y^{(1,1)})\}\cup\{(x^{\prime(1,1)},y^{\prime(1,1)})\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{(x^{(i,1)},y^{(i,1)}),\ldots,(x^{(i,m)},y^{(i,m)})\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:(x^{(1,j)},y^{(1,j)})=(0,0),\quad(x^{(1,1)},y^{(1,1)})=(bv,\frac{\beta}{b}),\\ &\qquad\qquad\qquad\text{ and }(x^{\prime(1,1)},y^{\prime(1,1)})=(0,0).\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:(x^{(i,j)},y^{(i,j)})=(0,0).\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:(x^{(i,j)},y^{(i,j)})=(v,-\alpha).\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:(x^{(i,j)},y^{(i,j)})=(bv,\frac{\beta}{b}).\end{split}$
Let $\gamma\leq\frac{1}{L}$ the learning rate. Following this setup, we first describe the form of the parameters and gradients throughout the training process, and then state the conditions under which we establish our instability result, assuming $\theta_{0}=\theta_{0}^{\prime}=0$. For $t\in\{0,\ldots,T-1\}$, and $\theta_{t}=\lambda_{t}v$, $\theta^{\prime}_{t}=\lambda^{\prime}_{t}v$, $\lambda_{t},\lambda^{\prime}_{t}\in[\lambda^{*}_{E}\vcentcolon=-\frac{\alpha}{L},\lambda^{*}_{F}\vcentcolon=\frac{\beta}{b^{2}L}]$:
$g^{(N)}_{t}=0,\quad\quad\quad$ $\|g^{(E)}_{t}\|_{2}>\psi\|g^{(F)}_{t}\|_{2},$ $g^{(F)}_{t}=-\underbrace{\left(\beta-\lambda_{t}Lb^{2}\right)}_{>0\text{ and }<\frac{C}{\sqrt{L}}}v$
$g^{(N)\prime}_{t}=0,\quad\quad\quad$ $g^{(E)\prime}_{t}=\underbrace{\left(\lambda^{\prime}_{t}L+\alpha\right)}_{>0\text{ and }<\frac{C}{\sqrt{L}}}v,$ $\|g^{(F)\prime}_{t}\|_{2}>\|g^{(E)\prime}_{t}\|_{2}$
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size nfn-f. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule, for all $t\in\{0,\ldots,T-1\}$:
$S^{*}_{t}\in\underset{S\subseteq\{1,\ldots,n\},|S|=n-f}{\text{argmin}}\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}\left(g^{(i)}_{t}-\overline{g}_{S}\right){\left(g^{(i)}_{t}-\overline{g}_{S}\right)}^{\intercal}\right).$
With a slight abuse of notation, we denote, for instance, the set $\{1\}\cup N\cup E$ by $\{1,N,E\}$. To uncover the source of instability, we consider the following two key conditions:
  • (C1)
    As shown in the $(C1)$ condition of Theorem D.2, the core tension in the choice of the $\mathrm{SMEA}$ aggregation rule ultimately reduces to a comparison between the sets $\{1,N,E\}$ and $\{1,N,F\}$, a consideration that also applies here. $S^{*\prime}_{0}=\{1,N,E\}\underset{\mathrm{SMEA}}{\Longleftrightarrow}\|g^{(F)\prime}_{0}\|_{2}=\beta\sqrt{L}>\alpha\sqrt{L}=\|g^{(E)\prime}_{0}\|_{2}$. Hence, $\theta^{\prime}_{1}=p_{\Theta}(-\gamma\frac{f}{n-f}g^{(E)}_{0})=p_{\Theta}(-\gamma\frac{f}{n-f}\alpha v)=0$. We then prove by induction, that for all $t\geq 0,S^{*\prime}_{t}=\{1,N,E\}$ and the unperturbed run remains at 0. The perturbed run remains at 0 until the differing samples are drawn.
  • (C2)
    Let denote
    $\lambda_{\max,t}^{\{1,N,E\}}\vcentcolon=\lambda_{\max}\left(\frac{1}{n-f}\sum_{i\in\{1,N,E\}}(g^{(i)}_{t}-\overline{g}_{S}){(g^{(i)}_{t}-\overline{g}_{S})}^{\intercal}\right),$
    and
    $\lambda_{\max,t}^{\{1,N,F\}}\vcentcolon=\lambda_{\max}\left(\frac{1}{n-f}\sum_{i\in\{1,N,F\}}(g^{(i)}_{t}-\overline{g}_{S}){(g^{(i)}_{t}-\overline{g}_{S})}^{\intercal}\right).$
    Assuming $T_{0}=t_{0}$ known, $S^{*}_{t_{0}-1}=\{1,N,F\}\underset{\mathrm{SMEA}\text{ and }(C1)}{\Longleftrightarrow}\frac{n-2(f+1)}{n-2}\|g^{(F)}_{t_{0}-1}\|_{2}<\|g^{(E)}_{t_{0}-1}\|_{2}$. In fact, by Lemma D.1 with $g_{A}=g^{(E)}_{t_{0}-1}$, $g_{B}=g^{(F)}_{t_{0}-1}$ and $g_{C}=0$:
    $S^{*}_{t_{0}-1}\in\underset{}{\text{argmin}}\bigg\{\lambda_{\max,t_{0}-1}^{\{1,N,F\}}=\frac{(f+1)(n-2f-1)}{{(n-f)}^{2}}\|g^{(F)}_{t_{0}-1}\|_{2}^{2},\\ \qquad\qquad\quad\lambda_{\max,t_{0}-1}^{\{1,N,E\}}=\frac{f(n-2f-1)}{{(n-f)}^{2}}\|g^{(E)}_{t_{0}-1}\|_{2}^{2}+\frac{(n-2f-1)}{{(n-f)}^{2}}\|g^{(F)}_{t_{0}-1}\|_{2}^{2}+\frac{f}{{(n-f)}^{2}}{\left(\|g^{(E)}_{t_{0}-1}\|_{2}+\|g^{(F)}_{t_{0}-1}\|_{2}\right)}^{2}\bigg\}.$
    Hence,
    $S^{*}_{t_{0}-1}=\{1,N,F\}$ $\Leftrightarrow\left(n-2f-1\right)(\|g^{(F)}_{t_{0}-1}\|_{2}^{2}-\|g^{(E)}_{t_{0}-1}\|_{2}^{2})<{(\|g^{(F)}_{t_{0}-1}\|_{2}+\|g^{(E)}_{t_{0}-1}\|_{2})}^{2}$
    $\Leftrightarrow\left(n-2f-1\right)(\|g^{(F)}_{t_{0}-1}\|_{2}-\|g^{(E)}_{t_{0}-1}\|_{2})<\|g^{(F)}_{t_{0}-1}\|_{2}+\|g^{(E)}_{t_{0}-1}\|_{2}$
    $\Leftrightarrow\psi\|g^{(F)}_{t_{0}-1}\|_{2}<\|g^{(E)}_{t_{0}-1}\|_{2}$
    which holds by construction of our initialization, $\psi\|g^{(F)}_{t_{0}-1}\|_{2}<\|g^{(E)}_{t_{0}-1}\|_{2}<\|g^{(F)}_{t_{0}-1}\|_{2}$.
    We then prove that for the next step, $S^{*}_{t_{0}}=\{1,N,F\}$ in any case. This is sufficient to prove by induction that for all $t\geq t_{0},S^{*}_{t}=\{1,N,F\}$ and the perturbed run converges to the minimizer of the subgroup $F$. In fact, for $t\geq t_{0}$, $\|g^{(F)}_{t}\|_{2}$ decreases and $\|g^{(E)}_{t}\|_{2}$ increases.
    • (1)
      with probability $\frac{1}{m}$, the differing samples are again drawn. As we already ensured the conditions in the previous step, we have $\psi\|g^{(F)}_{t_{0}}\|_{2}\leq\|g^{(F)}_{t_{0}-1}\|_{2}<g^{(E)}_{t_{0}-1}\|_{2}\leq\|g^{(E)}_{t_{0}}\|_{2}$, then $S^{*}_{t_{0}}=\{1,N,F\}$.
    • (2)
      with probability $1-\frac{1}{m}$ we have to ensure $S^{*}_{t_{0}}=\{1,N,F\}\underset{\mathrm{SMEA}}{\Longleftrightarrow}\|g^{(F)}_{t_{0}}\|_{2}\leq\|g^{(E)}_{t_{0}}\|_{2}$ with $\theta_{t_{0}}=-\gamma\frac{f+1}{n-f}g^{(F)}_{t_{0}-1}=\gamma\beta\frac{f+1}{n-f}v$ hence:
      $\|g^{(F)}_{t_{0}}\|_{2}=-\beta\left(1-\gamma\frac{f+1}{n-f}Lb^{2}\right)v,$
      $\|g^{(E)}_{t_{0}}\|_{2}=\max\{\beta,\beta\left(1-\epsilon+\gamma L\frac{f+1}{n-f}\right)\}v=\beta v$
    We proved that for $t\geq t_{0}-1$, $S^{*}_{t_{0}-1}=\{1,N,F\}$. Hence:
    $\mathbb{E}[\theta_{t}|T_{0}=t_{0}]=$ $\begin{cases}0&\text{if }t\leq t_{0}-1\\ \gamma\beta\frac{f+1}{n-f}v&\text{for }t=t_{0}\\ \mathbb{E}[\theta_{t-1}|T_{0}=t_{0}]-\gamma p\mathbb{E}[g^{(F)}_{t-1}|T_{0}=t_{0}]&\text{otherwise}.\end{cases}$
    with $\theta_{t}=\lambda_{t}v$, $p=\frac{f+\frac{1}{m}}{n-f}$, the recurrence for $t>t_{0}$ is given by:
    $\mathbb{E}[\lambda_{t}|T_{0}=t_{0}]$ $=\mathbb{E}[\lambda_{t-1}|T_{0}=t_{0}]+p\gamma(\beta-\mathbb{E}[\lambda_{t-1}|T_{0}=t_{0}]Lb^{2})$
    $=\left(1-p\gamma Lb^{2}\right)\mathbb{E}[\lambda_{t-1}|T_{0}=t_{0}]+p\gamma\beta$
    Hence:
    $\mathbb{E}[\theta_{t}|T_{0}=t_{0}]=$ $\begin{cases}0&\text{if }t\leq t_{0}-1\\ \frac{f+1}{f+\frac{1}{m}}\frac{\beta}{b^{2}L}\left(1-{(1-p\gamma Lb^{2})}^{t-t_{0}+1}\right)v&\text{otherwise}.\end{cases}$
Taking the total expectation, we obtain:
$\mathbb{E}[\theta_{T}]$ $=\sum_{t_{0}=1}^{T}\mathbb{E}[\theta_{T}|T_{0}=t_{0}]\mathbb{P}(T_{0}=t_{0})$
$=\frac{CT\frac{f+1}{f+\frac{1}{m}}}{Lm}\sum_{t_{0}=1}^{T}\left(1-{\left(1-\frac{p\gamma L}{T}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\frac{v}{\sqrt{L}}$
In the remainder of the proof, we bound $\mathbb{E}[\theta_{T}]$. Specifically, we leverage its analytical expression to establish a concise lower bound on uniform argument stability.
$\sum_{t_{0}=1}^{T}\left(1-{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}$ $\geq\sum_{t_{0}=1}^{T}\left(1-\frac{1}{1+p\gamma Lb^{2}(T-t_{0}+1)}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}$
$\geq\sum_{t_{0}=1}^{T}\frac{p\gamma L\frac{T-t_{0}+1}{T}}{1+p\gamma L\frac{T-t_{0}+1}{T}}{\left(1-\frac{1}{m}\right)}^{t_{0}-1}$
Where we used that ${\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\leq\frac{1}{1+p\gamma Lb^{2}(T-t_{0}+1)}$ for $T-t_{0}+1\geq 0$ and $1<-p\gamma Lb^{2}<0$. In fact using $\ln(1+x)\leq x$ for x>1x>-1,
$\ln\left[{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}(1+p\gamma Lb^{2}(T-t_{0}+1))\right]$ $=(T-t_{0}+1)\ln(1-p\gamma Lb^{2})+\ln(1+p\gamma Lb^{2}(T-t_{0}+1))$
$\leq-p\gamma Lb^{2}(T-t_{0}+1)+p\gamma Lb^{2}(T-t_{0}+1)=0.$
Also, as $\frac{1}{1+p\gamma L\frac{T-t_{0}+1}{T}}\geq\frac{1}{1+\gamma Lp}\geq\frac{1}{1+p}\geq\frac{1}{2}$, we have:
$\sum_{t_{0}=1}^{T}\left(1-{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\\ \qquad\qquad\geq\frac{p\gamma L}{2}\sum_{t_{0}=1}^{T}\left(1-\frac{t_{0}-1}{T}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)-\frac{p\gamma L}{2T}\sum_{t_{0}=1}^{T}(t_{0}-1){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)-\frac{p\gamma L}{2T}\left[m(m-1)-m^{2}{\left(1-\frac{1}{m}\right)}^{T}\left(1+\frac{T-1}{m}\right)\right]\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}\left[1-\frac{m-1}{T}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)\right]\hfill$
where we used the finite polylogarithmic sum formula in the second equality above.
If we additionally assume there exist a constant $\tau\geq c>0$ for an arbitrary c, such that $T\geq\tau m$, we then can prove that $K(T,m)=\left[1-\frac{m-1}{T}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)\right]\in\Omega(1)$. In fact this is trivial for $m=1$, then we prove that the quantity $K(T,m)$ increases with $T$, for $T\geq 1$ and $m\geq 2$. We prove by induction that $K(T+1)-K(T)=\frac{m-1}{T(T+1)}\left[1-{\left(1-\frac{1}{m}\right)}^{T}(1+\frac{T}{m})\right]>0$ which is equivalent to ${\left(1-\frac{1}{m}\right)}^{T}(1+\frac{T}{m})<1$. For $T=1$, $(1-\frac{1}{m})(1+\frac{1}{m})=1-\frac{1}{m^{2}}<1$. For $T+1>2$, $\frac{{\left(1-\frac{1}{m}\right)}^{T+1}(1+\frac{T+1}{m})}{{\left(1-\frac{1}{m}\right)}^{T}(1+\frac{T}{m})}=(1-\frac{1}{m})\frac{1+\frac{T+1}{m}}{1+\frac{T}{m}}<1$.
Hence:
$1-\frac{m-1}{T}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)$ $\geq 1-\frac{m-1}{\tau m}\left(1-{\left(1-\frac{1}{m}\right)}^{\tau m}\right)$
$\underset{(1)}{\geq}1-\frac{1}{\tau}(1-\frac{1}{m})\left(1-e^{-\tau}(1-\frac{\tau}{m})\right)$
$\underset{(2)}{\geq}1-\frac{1-e^{-\tau}}{\tau}$
$\geq 1-\frac{1-e^{-c}}{c}>0$
where we used in the second inequality $(1)$ that ${\left(1-\frac{1}{m}\right)}^{\tau m}\geq e^{-\tau}(1-\frac{\tau}{m})$ (we can prove it by taking the logarithm and compare their Taylor expansion). We also used in the third inequality $(2)$ that $1-\frac{1}{\tau}(1-\frac{1}{m})\left(1-e^{-\tau}(1-\frac{\tau}{m})\right)-1-\frac{1-e^{-\tau}}{\tau}=\frac{1}{\tau m}\left(1-(1+\tau)e^{-\tau}+\frac{\tau e^{-\tau}}{m}\right)\geq\frac{e^{-\tau}}{m^{2}}\geq 0$.
Wrapping-up:
$\mathbb{E}[\|\theta_{T}-\theta^{\prime}_{T}\|_{2}]\geq\|\mathbb{E}[\theta_{T}]\|_{2}\geq\frac{1}{2}(1-\frac{1-e^{-c}}{c})\frac{f+1}{f+\frac{1}{m}}p\gamma CT\in\Omega(p\gamma CT)$
Building on this lower bound for uniform argument stability, we then establish a connection to uniform stability, allowing us to complete the comparison with the upper bound from Theorem 3.3.
$\mathbb{E}\left[\ell(\theta_{T};(v,-\frac{C}{\sqrt{L}}))-\ell(\theta^{\prime}_{T};(v,-\frac{C}{\sqrt{L}}))\right]=C\sqrt{L}\mathbb{E}\lambda_{T}+\frac{C^{2}}{2L}-\frac{C^{2}}{2L}\in\Omega\left(pC^{2}\gamma T\right)\qed$
Appendix E DEFERED PROOFS AND ADDITIONAL DETAILS FROM SECTION 4
E.1 Proof of Lemmas 4.1 and 4.2
Below we compute the generalization error of a learning algorithm in the setting explicited in Section 4.
Lemma 4.1.
Consider the setting described in Section 2, with $m=1$, regardless of the assumed threat model. There exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, data distributions ${\{p_{i}\}}_{i\in\mathcal{H}}$ over $\mathcal{Z}$, such that for any distributed algorithm $\mathcal{A}$, we have:
$\left|\mathbb{E}_{\mathcal{A},\mathcal{S}\sim\otimes_{i\in\mathcal{H}}\left(p_{i}^{\otimes m}\right)}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]\right|=\frac{1}{4(n-f)}\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}\left[\ell(\mathcal{A}(S);z)-\ell(\mathcal{A}(S^{\prime});z)\right],$ (16)
where the dependence on the threat model arises solely through the stability term.
Proof.
We consider a linear loss function $z\in[-C,C],\theta\in\mathbb{R}\mapsto z\theta$, and we assume every honest local distribution is a Dirac (regardless of the value where the singularity is), except for a “pivot” honest worker (indexed $1$ without loss of generality) whose distribution is a mixture of Dirac $p_{1}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{-C}$. In this case, when the number of local samples is fixed to one ($m=1$), the two possible datasets are always neighboring. Denote them $S^{(-C)}$ and $S^{(0)}$. We denote the parameters resulting from $\mathcal{A}$ on the two different datasets by $\mathcal{A}(S^{(0)})=\theta_{T}^{(0)}$ and $\mathcal{A}(S^{(-C)})=\theta_{T}^{(-C)}$.
In this setup, we can compute the generalization error exactly as follows:
$\mathbb{E}_{\mathcal{S}}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]$ $=\frac{1}{|\mathcal{H}|}\mathbb{E}_{\mathcal{S}}\big[\mathbb{E}_{z\sim p_{1}}[\theta_{T}^{(z_{1})}z]-\theta_{T}^{(z_{1})}z_{1}\big]$
$=\frac{1}{|\mathcal{H}|}\mathbb{E}_{\mathcal{S}}\big[\frac{1}{2}\theta_{T}^{(z_{1})}\times 0+\frac{1}{2}\theta_{T}^{(z_{1})}\times(-C)-\theta_{T}^{(z_{1})}z_{1}\big]$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{C}{2}\mathbb{E}_{\mathcal{S}}\theta^{(z_{1})}_{T}-\mathbb{E}_{\mathcal{S}}\big[\theta_{T}^{(z_{1})}z_{1}\big]\right)$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{C}{2}\left(\frac{1}{2}\theta^{(0)}_{T}+\frac{1}{2}\theta^{(-C)}_{T}\right)-\frac{1}{2}\theta^{(0)}_{T}\times 0-\frac{1}{2}\theta^{(-C)}_{T}\times(-C)\right)$
$=\frac{C}{|\mathcal{H}|}\left(-\frac{1}{4}\theta^{(0)}_{T}-\frac{1}{4}\theta^{(-C)}_{T}+\frac{1}{2}\theta^{(-C)}_{T}\right)$
$=\frac{C\left(\theta^{(-C)}_{T}-\theta^{(0)}_{T}\right)}{4(n-f)}$ (17)
As the above computation is independent of the assumed threat model, we have,
$\left|\mathbb{E}_{\mathcal{A},\mathcal{S}}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]\right|=\frac{1}{4(n-f)}\sup_{z\in[-C,C]}\mathbb{E}_{\mathcal{A}}\left|\ell(\mathcal{A}(S^{(-C)});z)-\ell(\mathcal{A}(S^{(0)});z)\right|.\qed$
Building upon the theorem above, we leverage our stability analysis to demonstrate that the Byzantine failure and data poisoning threat models differ fundamentally in their capacity to affect the generalization error of distributed learning algorithms.
Theorem 4.2.
Consider the setting of Lemma 4.1, $\frac{n}{3}\leq f<\frac{n}{2}$, $\mathcal{A}=\mathrm{GD}$ with $\mathrm{SMEA}$. There exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, data distributions ${\{p_{i}\}}_{i\in\mathcal{H}}$ over $\mathcal{Z}$ and a Byzantine attack such that for any data poisoning attacks, we have
$\frac{\mathcal{E}^{\mathrm{byz}}_{gen}}{\mathcal{E}^{\mathrm{pois}}_{gen}}\in\Omega\left(\frac{n-f}{\sqrt{f(n-2f)}}\right),$ (18)
where $\mathcal{E}^{\mathrm{byz}}_{gen}$ and $\mathcal{E}^{\mathrm{pois}}_{gen}$ denote generalization errors under the Byzantine and data poisoning attacks, respectively.
Proof.
We specialize Lemma 4.1 to $\mathcal{A}=\mathrm{GD}$ with $\mathrm{SMEA}$. If we assume every honest local distribution is a Dirac with singularity at the value defined in the paragraph (i) of the proof of Theorem D.2—except for a “pivot” honest worker, indexed $1$, whose distribution is a mixture $p_{1}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{-C}$, then the only two possible datasets for honest workers align with the worst-case Byzantine and data-poisoning scenarios considered in our lower bounds (see Sections D.2.1 and C.3 for details of the datasets used). Hence, using the attacks described therein demonstrates that the generalization error exhibits the same intrinsic gap between the two threat models, consistent with the uniform stability analysis, as illustrated in Figure 1. That is $\mathcal{E}^{\mathrm{byz}}_{gen}$ scales according to eq. 4 and $\mathcal{E}^{\mathrm{pois}}_{gen}$ scales according to eq. 7 (both scaled by $\frac{1}{4\left(n-f\right)}$), as revealed by section E.1. ∎
E.2 Numerical analysis additional details
We describe here the details of our numerical experiment presented in Section 4 into two points: (i) the problem instantiation; (ii) our tailored Byzantine attack.
  • (i)
    We instantiate the setting from the proof of Theorem D.2 (see full details therein). That is, we consider $\mathcal{A}=\mathrm{GD}$ under a linear loss function, using the parameters $C=1$, $\gamma=1$, $T=5$, $n=15$, $m=1$. We make vary $f$ from $1$ to $7$, and for each value instantiate the dataset presented in the proof of Theorem D.2.
  • (ii)
    In our tailored Byzantine failures, the identities of the Byzantine workers remain fixed throughout the optimization process, but vary depending on the number of Byzantine workers $f$. The set of Byzantine identities is defined as $\mathcal{B}\vcentcolon=\{1,\ldots,n\}\setminus\mathcal{H}$, and is summarized in the table below. As we will see, we optimize the values sent by Byzantine workers so that they are selected while also biasing the gradient statistics—either toward positive or negative values—depending on an event that triggers them to switch their communicated value. To construct a maximally impactful scenario, we choose the identities of the Byzantine workers to maximize the minimal variance over all subsets of size nfn-f. This increases their biasing power, as it hinges on the smallest variance among these subsets. For instance, when $f=2$, the Byzantine workers can be hidden in the dominant subset $N$; however, when $f=6$, they must be distributed across both competing subgroups $E$ and $F$ to maintain influence.
    $f=1:\mathcal{B}=\{2\}$ $f=4:\mathcal{B}=\{2,3,4,5\}$
    $f=2:\mathcal{B}=\{2,3\}$ $f=5:\mathcal{B}=\{2,3,4,5,6\}$
    $f=3:\mathcal{B}=\{2,3,4\}$ $f=6:\mathcal{B}=\{7,8,9,10,11,12\}$
    $f=7:\mathcal{B}=\{6,7,8,9,10,11,12\}$
    For instance, if a single sample from worker $1$ (without loss of generality) is changed from 0 to C-C (defining a neighboring dataset), the server receives a different average $\frac{1}{m}C<0$ that triggers the Byzantine workers to amplify the positive or negative parameter direction by adaptively crafting updates that remain within the aggregation rule’s selection range. Specifically, their algorithm is as follows:
    • If $g_{0}^{(1)}\geq 0$, all Byzantine workers send the value $\alpha_{f}$, where $\alpha_{f}$ is numerically chosen to be as large as possible under the constraint that the $\mathrm{SMEA}$ aggregation rule continues to select the values sent by Byzantine workers in subsequent steps.
    • If $g_{0}^{(1)}<0$, all Byzantine workers send the value $\beta_{f}$, where $\beta_{f}$ is numerically chosen to be as large as possible under the constraint that the $\mathrm{SMEA}$ aggregation rule continues to select the values sent by Byzantine workers in subsequent steps.
    The following pseudocode summarizes the behavior of the Byzantine workers. Let $\alpha\in\mathbb{R}$ and
    $S^{*}_{\alpha}\in\underset{\begin{subarray}{c}|S|=n-f\end{subarray}}{\text{argmin}}\text{Var}\left({\{g_{k}\}}_{k\in S\cap\mathcal{H}}\cup{\{\alpha\}}_{k\in S\cap\mathcal{B}}\right).$
    if $g_{0}^{(1)}\geq 0$ then
     return $\alpha_{f}=\inf\bigg\{\alpha;i\in S^{*}_{\alpha}\bigg\}+\epsilon$# amplifies the initial direction
    else
     return $\beta_{f}=\sup\bigg\{\beta;i\in S^{*}_{\beta}\bigg\}-\epsilon$# amplifies the initial direction
    Algorithm 2 Byzantine worker $i\notin\mathcal{H}$ behavior, $\epsilon=10^{-3}$
Appendix F REFINED BOUNDS UNDER BOUNDED HETEROGENEITY AND BOUNDED VARIANCE ASSUMPTIONS
For clarity and simplicity, we chose to emphasize our results under the broad bounded gradient assumption (i.e. Lipschitz-continuity of the loss function) in the main text. In this section, we present results derived under refined assumptions—namely, bounded heterogeneity and bounded variance—in place of the more general bounded gradient condition. These assumptions are more commonly used in the optimization literature (as opposed to the generalization literature).
We formally define these assumptions below.
Assumption F.1 (Bounded heterogeneity).
There exists $G<\infty$ such that for all $\theta\in\Theta$: $\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|\nabla\widehat{R}_{i}(\theta)-\nabla\widehat{R}_{\mathcal{H}}(\theta)\|_{2}^{2}\leq G^{2}$.
Assumption F.2 (Bounded variance).
There exists $\sigma<\infty$ such that for each honest workers $i\in\mathcal{H}$ and all $\theta\in\Theta$: $\frac{1}{m}\sum_{x\in\mathcal{D}_{i}}\|\nabla_{\theta}\ell(\theta;x)-\nabla\widehat{R}_{i}(\theta)\|_{2}^{2}\leq\sigma^{2}$.
As shown in the following, our analysis techniques can yield tighter bounds when assuming bounded heterogeneity and bounded variance. However, we note that these bounds do not provide additional conceptual insights within the scope of our discussion. Indeed, within the uniform stability framework, if we assume only bounded gradients and nothing more, there exist distributed learning settings where $G$ is a constant multiple of the Lipschitz constant $C$. For instance, consider a scenario where half of the honest workers produce gradients with norm $C$, and the other half produce gradients with norm C-C; in this case, $G$ would be $C$.
As an example, we derive bounds on the expected spectral norm of the empirical covariance matrix of honest workers’ gradients that are expressed in terms of $G$ and $\sigma$. These estimates can be directly used in Theorem 3.1 to provide tighter bounds under additional assumptions of bounded heterogeneity and bounded variance.
Lemma F.1.
Consider the setting described in Section 2 under Byzantine attacks with Assumptions F.1 and F.2. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa)$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
$\mathbb{E}_{\mathcal{A}}\|\Sigma_{\mathcal{H},t}\|_{sp}=\mathbb{E}_{\mathcal{A}}[\lambda_{\max}(\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}(g^{(i)}_{t}-\overline{g}_{t}){(g^{(i)}_{t}-\overline{g}_{t})}^{\intercal})]\leq\sigma^{2}+G^{2}.$
Proof.
In the following, we derive a bound for the stochastic case $\mathcal{A}=\mathrm{SGD}$; the result naturally extends to the deterministic setting $\mathcal{A}=\mathrm{GD}$ as well. The proof of the refined upper bound leverages the ability to compute the expectation of the controllable noise—specifically, the noise resulting from random sample selection $J^{(i)}_{t}$, $i\in\mathcal{H}$, $t\in\{0,\ldots,T-1\}$—while conditioning on the uncontrollable noise introduced by Byzantine workers. Since the bound is independent of the latter, we establish the result by applying the law of total expectation.
Let $t\in\{0,\ldots,T-1\}$,
$\Delta_{t}=\lambda_{\max}(\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}(g^{(i)}_{t}-\overline{g}_{t}){(g^{(i)}_{t}-\overline{g}_{t})}^{\intercal})=\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\overline{g}_{t}\rangle^{2}.$
We decompose,
$g^{(i)}_{t}-\overline{g}_{t}=g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}$
so that we have:
$\langle v,g^{(i)}_{t}-\overline{g}_{t}\rangle^{2}=\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle^{2}+\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle^{2}\\ +2\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle.$
Upon averaging, taking the supremum over the unit ball, and then taking total expectations, we get:
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\overline{g}_{t}\rangle^{2}\right]\leq\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle^{2}\right]\\ +\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle^{2}\right]\\ +2\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle\right]$ (19)
We show that the last term in eq. 19 is non-positive. In fact, with $M=N+{N}^{\intercal}$,
$N=\sum_{i\in\mathcal{H}}(g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}){(\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t}))}^{\intercal}$
and using Lemma D.4 from Allouah et al. (2023b):
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}2\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle\right]\\ \quad=\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\langle v,Mv\rangle\right]\leq 9^{d}\sup_{\|v\|_{2}\leq 1}\mathbb{E}\left[\langle v,Mv\rangle\right]\hfill\\ \quad=9^{d}\sup_{\|v\|_{2}\leq 1}2\mathbb{E}\sum_{i\in\mathcal{H}}\langle v,\underbrace{\mathbb{E}[g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}|\theta_{t}]}_{=0}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle$
We now bound the second term in eq. 19:
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle^{2}\right]$ $\leq\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})\rangle^{2}\right]$
$\leq\mathbb{E}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\underbrace{\mathbb{E}_{J^{(i)}_{t}}[\|g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})\|^{2}_{2}]}_{\leq\sigma^{2}}\right]$
And finally we bound the first term in eq. 19 with F.1:
$\mathbb{E}\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle^{2}$ $\leq\sup_{\theta\in\Theta}\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,\nabla\widehat{R}_{i}(\theta)-\nabla\widehat{R}_{\mathcal{H}}(\theta)\rangle^{2}$
$=\sup_{\theta\in\Theta}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|\nabla\widehat{R}_{i}(\theta)-\nabla\widehat{R}_{\mathcal{H}}(\theta)\|^{2}_{2}\leq G^{2}\qed$
Appendix G HIGH-PROBABILITY EXCESS RISK BOUNDS FOR ROBUST DISTRIBUTED LEARNING
In this section, we establish high-probability excess risk bounds (see Equation 1) under smooth and strongly convex assumptions for robust distributed learning algorithms, providing insights into their generalization performance. Specifically, we combine a deterministic optimization error bound—adapted directly from prior work from Farhadkhani et al. (2024b)—with our high-probability stability bound.
High probability bound under uniform stability framework are derived as follows. The change of a single training sample has a small effect on a uniformly stable algorithm. This broadly implies low variance in the difference between empirical and population risk. If this difference also has low expectation, the empirical risk should accurately approximate the population risk (Bousquet and Elisseeff, 2002). In fact, we directly adapt the results from Bousquet et al. (2020) to our framework and state the following high-probability bound.
Lemma G.1.
Consider the setting described in Section 2, under either Byzantine failures or data poisoning. Let $\mathcal{A}$ an $\varepsilon$-uniformly stable. Assume $\forall z\in\mathcal{Z},\ell(\cdot,z)\leq\ell_{\infty}$, then we have that for any $0<\delta<1$, with probability at least $1-\delta$:
$|R_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))-\widehat{R}_{\mathcal{H}}(\mathcal{A}(\mathcal{S}))|\in\mathcal{O}\Bigg(\min\bigg\{\varepsilon\ln(|\mathcal{H}|m)\ln(\frac{1}{\delta})+\frac{\ell_{\infty}}{\sqrt{|\mathcal{H}|m}}\sqrt{\ln(\frac{1}{\delta})},\left(\sqrt{\varepsilon\ell_{\infty}}+\frac{\ell_{\infty}}{\sqrt{|\mathcal{H}|m}}\right)\sqrt{\ln(\frac{1}{\delta})}\bigg\}\Bigg)$
Data poisoning.
By adapting the optimization bound technique from Farhadkhani et al. (2024b, Theorem 4) to $\mathrm{GD}$ algorithm with $\mathrm{SMEA}$ under data poisoning, we show that for a constant number of steps—specifically, when the step count satisfies $T\geq|\ln(4\frac{f+1/m}{n-f}+\kappa)/\ln(1-\frac{\mu}{L})|\in\mathcal{O}(1)$—the expected excess risk is of the order of:
$\frac{C^{2}}{\mu}\mathcal{O}\left(\frac{f}{n-f}+\frac{1}{m(n-f)}\right).$
Additionally, we provide a high-probability bound on the excess risk using Lemma G.1. Specifically, for any $0<\delta<1$, with probability at least $1-\delta$ the excess risk lies within:
$\mathcal{O}\Bigg(\min\bigg\{\frac{f}{n-f}C^{2}+\left(\frac{f}{n-f}+\frac{1}{m(n-f)}\right)C^{2}\ln((n-f)m)\ln(\frac{1}{\delta})+\frac{\ell_{\infty}}{\sqrt{(n-f)m}}\sqrt{\ln(\frac{1}{\delta})},\\ \frac{f}{n-f}C^{2}+\left(C\sqrt{\ell_{\infty}\left(\frac{f}{n-f}+\frac{1}{m(n-f)}\right)}+\frac{\ell_{\infty}}{\sqrt{(n-f)m}}\right)\sqrt{\ln(\frac{1}{\delta})}\bigg\}\Bigg)$
In the absence of a minimax statistical lower bound for the heterogeneous case—and acknowledging that the comparison is imperfect due to differing assumptions—it remains informative to compare our resulting bound with those from Yin et al. (2018); Zhu et al. (2023), as well as Allouah et al. (2023a, Proposition 1).
Byzantine failures.
Under Byzantine failures, the derivation proceeds similarly and yields an analogous bound, with a dependence on $\sqrt{\kappa}$ replacing the $\frac{f}{n-f}$ term in the uniform stability bound.
Appendix H ADDITIONAL DISCUSSIONS
Why the difference in threat models (Byzantine failures versus data poisoning) affects generalization but not optimization?
The key intuition behind why the difference in threat models affects generalization but not optimization is as follows. Optimization error analysis focuses on how accurately the true descent direction is estimated at each step, whereas generalization error analysis examines the algorithm’s sensitivity to individual training samples, independent of how informative the estimated direction is for optimization. In the worst case, both types of corruption similarly impair the estimation of the optimization direction. However, stability analysis benefits from the additional regularity present in data poisoning, where corrupted gradients still correspond to gradients of a smooth loss function (i.e., they remain Lipschitz)—unlike Byzantine failures, which allow for arbitrary, potentially non-smooth corruptions.
Specifically, optimization error analysis relies on the smoothness of the empirical loss over honest workers (e.g., Theorem 1 in Allouah et al., 2023a) to control its decrease along the optimization trajectory, typically via terms like $\widehat{R}_{\mathcal{H}}(\theta_{t+1})-\widehat{R}_{\mathcal{H}}(\theta_{t})$ or $\|\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\|^{2}_{2}$. This smoothness-based control applies equally to both threat models, as the Lipschitzness of corrupted gradients is not invoked, and yields optimal bounds. In contrast, stability analysis aims to bound the difference $\ell(\theta_{t},z)-\ell(\theta^{\prime}_{t},z)$ between the losses along two optimization trajectories $\theta_{t}$ and $\theta^{\prime}_{t}$ produced by neighboring datasets. Here, smoothness can be exploited under data poisoning—where corrupted updates remain gradients of a smooth loss function (cf. proof of Theorem 3.3)—but not under Byzantine failures, which require an intermediate comparison step (cf. proof sketch in Section C.2).
Furthermore, prior work by Allouah et al. (2023a) have shown that for the optimization error, $(f,\kappa)$-robustness with $\kappa\in\mathcal{O}(f/n)$ suffices to achieve a lower bound obtained with poisoned data. A matching upper-bound is obtained for both Byzantine failures and data poisoning threat models (Allouah et al., 2023a, Theorem 1 & Proposition 1), meaning that we cannot hope to improve it, even using the Lipschitzness of corrupted gradients under data poisoning.
On the relevance of worst-case bounds to practical learning scenarios.
Although variations in attack strength and implementation may contribute to the observed gap in generalization performance between the two threat models, our results indicate that this gap cannot be explained solely by the fact that data poisoning attacks used in practice are often empirically weak (suboptimal within their threat model), whereas Byzantine attacks used in practice tend to be empirically strong (near-optimal). Rather, the gap also reflects an inherent difference in harm capacity (Section 4).
We recall that a gap in test accuracy has been consistently reported in prior work on realistic datasets, further motivating our investigation. For example, (Allouah et al., 2023a, Table 2) compares Byzantine attacks such as $\mathrm{ALIE}$, $\mathrm{FOE}$, and $\mathrm{SF}$ with data poisoning attacks such as $\mathrm{LF}$ and $\mathrm{Mimic}$, showing a clear disparity in their impact.
Does our conclusions would fundamentally change for other popular first-order optimization methods (e.g. momentum, variance reduction, multiple local steps)?
We believe our proof techniques are broadly applicable to robust first-order optimization, so the fundamental results should remain consistent. Quantitatively characterizing this phenomenon for specific algorithms is an interesting direction for future work. More concretely, similar recursions can be derived for stochastic heavy-ball momentum, using the proof techniques of Section C.2 (under Byzantine failures), and the refined approach in the proof of Theorem 3.3 (under data poisoning). These derivations reveal a similar gap as in robust distributed $\mathrm{GD}$ and $\mathrm{SGD}$: a $\sqrt{\kappa}$ dependency appears under Byzantine failures, while the tighter bound $\frac{f}{n-f}$ arises under data poisoning. However, fully analyzing these more complex algorithms introduces additional technical challenges. To maintain clarity and focus, we chose to highlight the $\mathrm{GD}$$\mathrm{SGD}$ case, which already illustrates the core insights.
Many practical algorithms perform multiple local $\mathrm{SGD}$ steps to reduce communication overhead (e.g., $\mathrm{FedAvg}$). However, our focus in this work is not on practical communication efficiency, but rather on understanding the generalization behavior in the presence of misbehaving agents in the most direct way for simplicity and clarity. We thus chose to analyze the setting where each client performs a single gradient step which provides a clearer and more focused framework to present and understand the fundamental algorithmic stability properties. We believe our insights extend to the multi-step case, and see no reason why the general conclusion of our paper would be impacted. Exploring this formally is an interesting direction for future work.
Discussion on other threat model formulations.
Other threat models could be investigated in future work, building upon the proof techniques developed in our work. To offer some preliminary context, we briefly discuss two relevant threat model formulations below.
1. Locally-poisonous threat model. The recent work by Farhadkhani et al. (2024b), which studies the locally-poisonous threat model, indicates that this threat model tends to have a fundamentally lower impact on optimization error compared to the Byzantine failures and data poisoning threat models considered in our paper. Studying stability and generalization under locally-poisonous threat model is an interesting future work.
2. Backdoor (or targeted) attacks. While important, this threat model differs in goals and evaluation from our study of generalization error under untargeted attacks. That said, they can be seen as special cases of Byzantine failures and data poisoning. Extending our theory to handle targeted attacks like backdoors remains an interesting direction for future work.